### v1

# Splitting of uniform bundles on quadrics

Abstract.
We show that there exist only constant morphisms from $\mathbb{Q}^{2n+1}(n\geq 1)$ to $\mathbb{G}(l,2n+1)$ if $l$ is even $(0<l<2n)$ and $(l,2n+1)$ is not $(2,5)$. As an application, we prove on $\mathbb{Q}^{2m+1}$ and $\mathbb{Q}^{2m+2}(m\geq 3)$, any uniform bundle of rank $2m$ splits.
1.Xinyi Fang is supported by innovation Action Plan (Basic research projects) of Science and Technology Commission of Shanghai Municipality (Grant No.21JC1401900)
2.Duo Li is supported by National Natural Science Foundation of China (Grant No. 12001547)
Keywords: uniform bundle; quadric; splitting of vector bundles.
MSC: 14M15; 14M17; 14J60.
1. Introduction
In this article, we assume all the varieties are defined over $\mathbb{C}$. Assume that $X$ is a rational homogeneous variety of Picard number $1$ (we call $X$ a generalised Grassmannian for short), then $X$ is swepted by lines. Let $E$ be a vector bundle on $X$. We consider its restriction $E|_{L}(\simeq\mathcal{O}(a_{1})\oplus\cdots\oplus\mathcal{O}(a_{k}))$ to any line $L\subseteq X$. If the splitting type $(a_{1},\dots,a_{k})$ of $E|_{L}$ is independent of the choice of $L$, we call $E$ a uniform bundle.
Every uniform bundle on $\mathbb{P}^{n}$ whose rank is smaller than $n$ splits (cf. the main theorems of [2] and [8]). On a Grassmannian $Gr(k,n+1)(k\leq n+1-k)$, any uniform bundle whose rank is smaller than $k$ splits (cf. [4, Theorem 1]). On some other generalised Grassmannians, there are similar results, see [1, Theorem 1.5], [5, Theorem 3.1] and [7, Theorem 5.6].
In a recent article [3], for an arbitrary generalised Grassmannian $X$, we prove that once the rank of a uniform bundle $E$ is smaller than $e.d.(\mathrm{VMRT})+1$ (for the definition of $e.d.(\mathrm{VMRT})$, see [3, Section $3$]), the bundle $E$ necessarily splits. Moreover, for most generalised Grassmannians, we classify unsplit uniform bundles of minimal ranks. The upper bound $e.d.(\mathrm{VMRT})$ is optimal for many generalised Grassmannians, for example it is optimal for $\mathbb{Q}^{n}(3\leq n\leq 6)$ (see [3, Theorem 1.1 (2)]). However, we will see that this upper bound is not optimal for higher dimensional quadrics.
In this article, we first show that there exist only constant morphisms from $\mathbb{Q}^{2n+1}(n\geq 1)$ to $\mathbb{G}(l,2n+1)$ if $l$ is even $(0<l<2n)$ and $(l,2n+1)$ is not $(2,5)$. Combining this result with the method of [2], we prove that every uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}(n\geq 3)$ splits and the upper bound $2n$ is optimal for $\mathbb{Q}^{2n+1}$. This gives the first known example such that the optimal upper bound is bigger than $e.d.(\mathrm{VMRT})$.
This article is organised as follows: in Section $2,$ we study morphisms from $\mathbb{Q}^{2n+1}(n\geq 1)$ to $\mathbb{G}(l,2n+1)$ and the main result of this section is Proposition 2.1; in Section $3,$ we study uniform bundles on smooth quadrics and the main result of this section is Theorem 3.11. Furthermore, we also study uniform bundles on $B_{n}/P_{k}$ and $D_{n}/P_{k}$, see Corollary 3.13, Corollary 3.14, Corollary 3.15 and Corollary 3.16.
2. Morphisms from quadrics to Grassmannians
Proposition 2.1.
There exist only constant morphisms from $\mathbb{Q}^{2n+1}(n\geq 1)$ to $\mathbb{G}(l,2n+1)$ if $l$ is even $(0<l<2n)$ and $(l,2n+1)$ is not $(2,5)$.
Proof.
We mainly follow the proof of the main theorem in [10]. Let $H$ be the cohomology class of a hyperplane on $\mathbb{Q}^{2n+1}$. The cohomology ring of $\mathbb{Q}^{2n+1}$ is
$H^{\bullet}(\mathbb{Q}^{2n+1},\mathbb{Z})=\mathbb{Z}\oplus\mathbb{Z}H\oplus\cdots\oplus\mathbb{Z}H^{n-1}\oplus\mathbb{Z}\frac{H^{n}}{2}\oplus\cdots\oplus\mathbb{Z}\frac{H^{2n+1}}{2}.$
Let $U$ (resp. $Q$) be the universal subbundle (resp. quotient bundle) on $\mathbb{G}(l,2n+1)$. Suppose that $f:\mathbb{Q}^{2n+1}\rightarrow\mathbb{G}(l,2n+1)$ is a non-constant morphism. Let $c_{i}$ and $d_{j}$ be rational numbers satisfying
$c_{i}(f^{*}U^{\vee})=c_{i}H^{i}\text{ for }1\leq i\leq l+1~{}\text{ and }~{}c_{j}(f^{*}Q)=d_{j}H^{j}\text{ for }1\leq j\leq 2n+1-l.$
By [9, Proposition 2.1 (i)], both $c_{i}(1\leq i\leq l+1)$ and $d_{j}(1\leq j\leq 2n+1-l)$ are non-negative. We note that for any $1\leq i,j<n$, $c_{i}$ and $d_{j}$ are integers. For any $i,j\geq n$, $2c_{i}$ and $2d_{j}$ are integers. Then from the exact sequence $0\rightarrow f^{*}U\rightarrow\mathcal{O}_{\mathbb{Q}^{2n+1}}^{\oplus 2n+2}\rightarrow f^{*}Q\rightarrow 0$, we get the equality of polynomials:
$(1-c_{1}t+c_{2}t^{2}+\cdots+(-1)^{l+1}c_{l+1}t^{l+1})(1+d_{1}t+\cdots+d_{2n+1-l}t^{2n+1-l})$ (2.1)
$=$ $1+(-1)^{l+1}c_{l+1}d_{2n+1-l}t^{2n+2}.$
If $c_{l+1}d_{2n+1-l}$ is 00, then by (2.1) both $c_{1}$ and $d_{1}$ are zero, which implies that $f$ is constant. So we may assume the numbers $c_{1},d_{1},c_{l+1}$ and $d_{2n+1-l}$ are non-zero. By [9, Proposition 2.1 (ii)], all the rational numbers $c_{1},c_{2},\dots,c_{l+1}$ and $d_{1},d_{2},\dots,d_{2n+1-l}$ are positive.
Let $a$ be $2n+2]{c_{l+1}d_{2n+1-l}}$. We set $C_{i}:=\frac{c_{i}}{a^{i}}(1\leq i\leq l+1)$ and $D_{j}:=\frac{d_{j}}{a^{j}}(1\leq j\leq 2n+1-l)$. We note that if $a,C_{i}$ and $D_{j}$ are all positive integers, then by the same proof of [10, Case (ii)], we can get a contradiction. Then it suffices to show that $a,C_{i}$ and $D_{j}$ are integers.
We firstly show that $a$ is an integer. Since $l$ is even, similar to the proof in [9, Lemma 3.3 (ii) and Case 1 of (iii)], we can show that $a$ is $\frac{c_{m+1}}{c_{m}}$, where $m$ is $\frac{l}{2}$. So $a$ is rational, we may assume $a=s/t$, where $s$ and $t$ are coprime positive integers. By definition, we have $a^{2n+2}=c_{l+1}d_{2n+1-l}$. Since $2c_{l+1}$ and $2d_{2n+1-l}$ are integers, $4a^{2n+2}$ is an integer. So $t^{2n+2}$ divides $4$, which implies that $t$ is $1$, as $n$ is at least $1$. Hence $a$ is an integer.
We now show that $C_{i}(=\frac{c_{i}}{a^{i}})$ and $D_{j}(=\frac{d_{j}}{a^{j}})$ are integers. Let $F_{1}(x)\cdots F_{k}(x)$ be the irreducible factorization of $1-x^{2n+2}$ over $\mathbb{Z}[x]$ with $F_{l}(0)=1(1\leq l\leq k)$. Then $F_{l}(x)$ is also irreducible over $\mathbb{Q}[x]$ by Gauss Lemma. Then
$(1-c_{1}t+c_{2}t^{2}+\cdots+(-1)^{l+1}c_{l+1}t^{l+1})(1+d_{1}t+\cdots+d_{2n+1-l}t^{2n+1-l})$
$=$ $1+(-1)^{l+1}c_{l+1}d_{2n+1-l}t^{2n+2}=1-a^{2n+2}t^{2n+2}=F_{1}(at)\cdots F_{k}(at).$
Note that $F_{i}(at)$ is also irreducible over $\mathbb{Q}[t]$. We have
$1-c_{1}t+c_{2}t^{2}+\cdots+(-1)^{l+1}c_{l+1}t^{l+1}=F_{i_{1}}(at)\cdots F_{i_{k_{1}}}(at)~{}\text{and}~{}$
$1+d_{1}t+\cdots+d_{2n+1-l}t^{2n+1-l}=F_{j_{1}}(at)\cdots F_{j_{k_{2}}}(at).$
Since the coefficients of $F_{l}(x)$ are integers, $C_{i}(=\frac{c_{i}}{a^{i}})$ and $D_{j}(=\frac{d_{j}}{a^{j}})$ are integers. ∎
3. Uniform bundles on $\mathbb{Q}^{2n+1}(n\geq 3)$
We are going to use the method in [2] to show that any uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}(n\geq 3)$ splits. We firstly fix some notations.
Let $E$ be a uniform bundle on $X(=\mathbb{Q}^{2n+1})(n\geq 3)$ of rank $2n$. Assume that $E$ does not split, by [3, Proposition $3.7$], we may assume that the splitting type of $E$ is
$(\underbrace{0,\dots,0}_{l+1},\underbrace{-1,\dots,-1}_{2n-l-1})\quad(l+1\geq 2n-l-1\geq 1).$
We denote the moduli of lines on $X$ and the corresponding universal family by:
${\mathcal{U}(=B_{n+1}/P_{1,2})}$${\mathcal{M}(=B_{n+1}/P_{2})}$${X(=B_{n+1}/P_{1}).}$$\scriptstyle{q}$$\scriptstyle{p}$
where $\mathcal{M}$ is the moduli of lines and $\mathcal{U}$ is the universal family. The relative Harder-Narasimhan (H-N) filtration of $p^{*}E$ induces an exact sequence:
$0\rightarrow E_{1}(=q^{*}G_{1})\rightarrow p^{*}E\rightarrow E_{2}(=q^{*}G_{2}\otimes p^{*}\operatorname{\mathcal{O}}_{X}(-1))\rightarrow 0,$ (3.1)
where $G_{1}$ (resp. $G_{2}$) is a vector bundle on $\mathcal{M}$ of rank $l+1$ (resp. $12n-l-$ 1). For each $x\in X$, the restriction of relative H-N filtration to $p^{-1}(x)$ induces a morphism
$\psi_{x}:p^{-1}(x)(\cong\mathbb{Q}^{2n-1})\rightarrow Gr(l+1,2n)(\cong\mathbb{G}(l,2n-1)).$ (3.2)
By [3, Lemma $2.1$], we have the following description of cohomology rings:
$H^{\bullet}(X,\mathbb{Q})=\mathbb{Q}[X_{1}]/(X_{1}^{2n+2}),$
$H^{\bullet}(\mathcal{U},\mathbb{Q})=\mathbb{Q}[X_{1},X_{2}]/(\Sigma_{i}(X_{1}^{2},X_{2}^{2})_{n\leq i\leq n+1}),$
$H^{\bullet}(\mathcal{M},\mathbb{Q})=\mathbb{Q}[X_{1}+X_{2},X_{1}X_{2}]/(\Sigma_{i}(X_{1}^{2},X_{2}^{2})_{n\leq i\leq n+1}).$
For a bundle $F$ of rank $r$, the Chern polynomial of $F$ is defined as
$C_{F}(T):=T^{r}-c_{1}(F)T^{r-1}+\cdots+(-1)^{r}c_{r}(F).$
Let $E(T,X_{1})=\sum_{k=0}^{2n}e_{k}X_{1}^{k}T^{2n-k}(\in\mathbb{Q}[X_{1},T])$ and $S_{i}(T,X_{1},X_{2})(\in\mathbb{Q}[X_{1}+X_{2},X_{1}X_{2},T])(i=1,2)$ be homogeneous polynomials representing $C_{p^{*}E}(T)$ and $C_{q^{*}G_{i}}(T)$ in the cohomology rings respectively. There are equations
$E(T,X_{1})=C_{p^{*}E}(T)~{}\text{and}~{}S_{i}(T,X_{1},X_{2})=C_{q^{*}G_{i}}(T)(i=1,2).$
Let $R(X_{1},X_{2})$ be the polynomial $\Sigma_{n}(X_{1}^{2},X_{2}^{2})$. By (3.1), we have an equation of Chern polynomials:
$E(T,X_{1})-aR(X_{1},X_{2})=S_{1}(T,X_{1},X_{2})S_{2}(T+X_{1},X_{1},X_{2}).$ (*)
If $a$ is 00, then both $S_{1}(T,X_{1},X_{2})$ and $S_{2}(T+X_{1},X_{1},X_{2})$ are polynomials only in variables $T$ and $X_{1}$. Since $S_{i}(T,X_{1},X_{2})(i=1,2)$ are symmetrical in $X_{1}$ and $X_{2}$, we must have $S_{1}(T,X_{1},X_{2})=T^{l+1}$ and $S_{2}(T,X_{1},X_{2})=T^{2n-l-1}$. So $c_{1}(E_{1})$ and $c_{1}(E_{2})$ are 00, and $\psi_{x}$ is constant for each $x\in X$, which implies that $E$ splits (see, for example, [3, Proposition $3.5$]). Therefore, we have $a\neq 0$.
3.1. Approximate solutions
To solve the equation (*3), we use the concept of approximate solutions introduced in [2, Section 5]. The following definitions and propositions are basically from [2, Section 5] and the proofs are similar.
Definition 3.1.
A non-zero homogeneous polynomial $P(T,X_{1})$ with rational coefficients in variables $T$ and $X_{1}$ of degree $2n$ is called an approximate solution if $P(T,X_{1})-R(X_{1},X_{2})$ has a proper divisor $S(T,X_{1},X_{2})$ which is symmetrical in $X_{1}$ and $X_{2}$. We call such a divisor a symmetrical divisor.
Example 3.2.
By the the equation (*3), both $\frac{1}{a}E(T,X_{1})$ and $\frac{1}{a}E(T-X_{1},X_{1})$ are approximate solutions. We call them approximate solutions associated with (*3).
In the following lemma, there are some restrictions on the coefficient of $X_{1}^{2n}$ for an arbitrary approximate solution $P(T,X_{1})$.
Lemma 3.3.
Let $P(T,X_{1})=\sum_{k=0}^{2n}p_{k}X_{1}^{k}T^{2n-k}$ be an approximate solution. Then one of the followings holds.
$(1)$ Any symmetrical divisor of $P(T,X_{1})-R(X_{1},X_{2})$ is of degree one.
$(2)$ The coefficient $p_{2n}$ is 00 and the zero set of $P(0,1)-R(1,z)$ is ($(\{z|z^{2n+2}-1=0\})\backslash\{1,-1\}$.
$(3)$ The coefficient $p_{2n}$ is $1$ and the zero set of $P(0,1)-R(1,z)$ is ($(\{z|z^{2n}-1=0\}\cup\{0\})\backslash\{1,-1\}$.
Proof.
Let $S(T,X_{1},X_{2})$ be a symmetrical divisor of $P(T,X_{1})-R(X_{1},X_{2})$. Then $S(0,X_{1},X_{2})$ divides $p_{2n}X_{1}^{2n}-R(X_{1},X_{2})$. As $S(T,X_{1},X_{2})$ is symmetrical in $X_{1}$ and $X_{2}$, $S(0,X_{1},X_{2})$ divides $p_{2n}X_{2}^{2n}-R(X_{1},X_{2}).$ Therefore $S(0,X_{1},X_{2})$ divides $p_{2n}(X_{1}^{2n}-X_{2}^{2n})$. By the equation $R(X_{1},X_{2})(X_{1}^{2}-X_{2}^{2})=X_{1}^{2n+2}-X_{2}^{2n+2}$, we have
$(X_{1}^{2}-X_{2}^{2})(p_{2n}X_{1}^{2n}-R(X_{1},X_{2}))+X_{2}^{2}(X_{1}^{2n}-X_{2}^{2n})=(p_{2n}-1)X_{1}^{2n}(X_{1}^{2}-X_{2}^{2}).$
If $p_{2n}$ is not 00 and $1$, we have $S(0,X_{1},X_{2})\mid X_{1}^{2n}-X_{2}^{2n}$ and $S(0,X_{1},X_{2})\mid X_{1}^{2n}(X_{1}-X_{2})(X_{1}+X_{2})$. Since $S(0,X_{1},X_{2})$ is symmetrical in $X_{1}$ and $X_{2}$, $S(0,X_{1},X_{2})$ is $c(X_{1}+X_{2})$ for some $c\in\mathbb{Q}$. In particular, $\deg(S)$ is $1$.
If $p_{2n}$ is 00, then $P(0,1)-R(1,z)$ is $-\frac{z^{2n+2}-1}{z^{2}-1}(=-(z^{2n}+z^{2n-2}+\cdots+1))$. If $p_{2n}$ is $1$, then $P(0,1)-R(1,z)$ is $-z^{2}\frac{z^{2n}-1}{z^{2}-1}=(-(z^{2n}+z^{2n-2}+\cdots+z^{2}))$. Then the assertions follow immediately. ∎
Definition 3.4.
We call an approximate solution $P(T,X_{1})=\sum_{k=0}^{2n}p_{k}X_{1}^{k}T^{2n-k}$ a primitive approximate solution if $p_{2n}\in\{0,1\}$ and $P(T,X_{1})-R(X_{1},X_{2})$ has a symmetrical divisor $S_{0}(T,X_{1},X_{2})$ such that there is a $2(n-p_{2n}+1)$-th primitive unit root $y_{0}$ satisfying $S_{0}(0,1,y_{0})=0$.
As in [2], we have the following classifications of primitive approximate solutions.
Proposition 3.5.
Let $P(T,X_{1})$ be a primitive approximate solutions. If $p_{2n}$ is 00, we have $P(T,X_{1})=bT^{2n}$ for some $b\in\mathbb{Q}$. If $p_{2n}$ is $1$, we have $P(T,X_{1})=\Sigma_{n}(bT^{2},X_{1}^{2})$ for some $b\in\mathbb{Q}$.
The proof of Proposition 3.5 is similar to that of [2, Proposition 6.1 and Proposition 6.2], we leave them in the Appendix (see Propositions A.1 and A.2).
3.2. The case $l$ is even
We firstly aim to show that $l$ is not even and we will prove it by contradiction. Now suppose that $l$ is even. If $l$ is smaller than $22n-$ 2 and $(l,2n-1)$ is not $(2,5)$, the morphism $\psi_{x}$ (for the definition of $\psi_{x}$, see (3.2)) is constant for each $x\in X$ according to Proposition 2.1. Then $E$ splits. We exclude the remaining cases $l=2n-2$ and $(l,2n-1)=(2,5)$ by calculations.
Proposition 3.6.
There does not exist an unsplit uniform bundle of rank $2n$ whose splitting type is $(0,\dots,0,-1)$ on $\mathbb{Q}^{2n+1}(n\geq 3)$.
Proof.
Suppose $E$ is unsplit, by the same calculation as in [3, Theorem 4.3, Case I], we have $c_{1}(E_{2})=(X_{1}+X_{2})-X_{1}=X_{2}$. Let $f(X_{1},X_{2})$ be a homogeneous polynomial of degree $12n-$ 1 which is symmetrical in $X_{1}$ and $X_{2}$ and represents $c_{2n-1}(E_{1})$. By comparing the coefficients of $T^{0}$ on the left and right sides of the equation (*3), we get
$f(X_{1},X_{2})X_{2}=e_{2n}X_{1}^{2n}-a(X_{1}^{2n}+X_{1}^{2n-2}X_{2}^{2}+\cdots+X_{2}^{2n}).$
Then we must have $e_{2n}=a$ and $f(X_{1},X_{2})=-a(X_{1}^{2n-2}X_{2}+X_{1}^{2n-4}X_{2}^{3}+\cdots+X_{2}^{2n-1})$, contradicting to the assumption that $f$ is symmetrical in $X_{1}$ and $X_{2}$. ∎
We now exclude the case $(l,2n-1)=(2,5)$.
Proposition 3.7.
There does not exist an unsplit uniform of rank $6$ whose splitting type is $(0,0,0,-1,-1,-1)$ on $\mathbb{Q}^{7}$.
Proof.
Suppose $E$ is unsplit. Then $a$ in the equation (*3) is not 00 and $\frac{1}{a}E(T,X_{1})=\frac{1}{a}\sum_{k=0}^{6}e_{k}X_{1}^{k}T^{6-k}$ is an approximate solution which has a symmetrical divisor of degree $3$. By Lemma 3.3, we have $\frac{1}{a}e_{6}=0$ or $\frac{1}{a}e_{6}=1$. Let $f(X_{1},X_{2})$ be a homogeneous polynomial representing $c_{3}(E_{1})$.
If $\frac{1}{a}e_{6}$ is 00, we have $f(X_{1},X_{2})\mid R(X_{1},X_{2})(=X_{1}^{6}+X_{1}^{4}X_{2}^{2}+X_{1}^{2}X_{2}^{4}+X_{2}^{6})$. Since the prime factorization of $X_{1}^{6}+X_{1}^{4}X_{2}^{2}+X_{1}^{2}X_{2}^{4}+X_{2}^{6}$ over $\mathbb{Q}[X_{1},X_{2}]$ is $(X_{1}^{4}+X_{2}^{4})(X_{1}^{2}+X_{2}^{2})$, $R(X_{1},X_{2})$ has no divisor symmetrical in $X_{1}$ and $X_{2}$ of degree $3$.
If $\frac{1}{a}e_{6}$ is $1$, then $f(X_{1},X_{2})$ divides $X_{1}^{6}-R(X_{1},X_{2})(=-X_{2}^{2}(X_{1}^{4}+X_{1}^{2}X_{2}^{2}+X_{2}^{4}))$. The prime factorization of $X_{1}^{4}+X_{1}^{2}X_{2}^{2}+X_{2}^{4}$ over $\mathbb{Q}[X_{1},X_{2}]$ is $(X_{1}^{2}+X_{1}X_{2}+X_{2}^{2})(X_{1}^{2}-X_{1}X_{2}+X_{2}^{2})$. Therefore, $X_{1}^{6}-R(X_{1},X_{2})$ has no divisor symmetrical in $X_{1}$ and $X_{2}$ of degree $3$.
In both cases, we get contradictions. ∎
3.3. The case $l$ is odd
Suppose that $l$ is odd. We begin with a lemma.
Lemma 3.8.
When $l$ is odd, the equation $E(t,1)-aR(1,0)=0$ has no roots in $\mathbb{R}$.
Proof.
In the equation (*3): $E(T,X_{1})-aR(X_{1},X_{2})=S_{1}(T,X_{1},X_{2})S_{2}(T+X_{1},X_{1},X_{2})$, we let $X_{1}$ be 00 and let $X_{2}$ be $1$. Then we get an equation $T^{2n}-a=S_{1}(T,0,1)S_{2}(T,0,1)$. We write $S_{1}(T,0,X_{2})$ and $S_{2}(T,0,X_{2})$ as follows: $S_{1}(T,0,X_{2})=\sum_{i=0}^{l+1}a_{i}X_{2}^{i}T^{l+1-i}$ and $S_{2}(T,0,X_{2})=\sum_{j=0}^{2n-1-l}b_{j}X_{2}^{j}T^{2n-1-l-j}$, where $a_{i}(0\leq i\leq l+1)$ and $b_{j}(0\leq j\leq 2n-1-l)$ are rational numbers. Then $-a$ is $a_{l+1}b_{2n-1-l}$. We now wish to show $-a>0$. To this end, for any $x$ in $X$, we consider the embedding $i_{x}:p^{-1}(x)(=\mathbb{Q}^{2n-1})\hookrightarrow\mathcal{U}(=B_{n+1}/P_{1,2})$, which induces a morphism:
$i_{x}^{*}:H^{\bullet}(\mathcal{U},\mathbb{Q})\cong\mathbb{Q}[X_{1},X_{2}]/(\Sigma_{n}(X_{1}^{2},X_{2}^{2}),\Sigma_{n+1}(X_{1}^{2},X_{2}^{2}))\rightarrow\mathbb{Q}[X_{2}]/X_{2}^{2n}\cong H^{\bullet}(\mathbb{Q}^{2n-1},\mathbb{Q}).$
Under the above identifications, we have $S_{1}(T,0,X_{2})=C_{E_{1}|_{p^{-1}(x)}}(T)=C_{\psi_{x}^{*}U}(T)$ and $S_{2}(T,0,X_{2})=C_{E_{2}|_{p^{-1}(x)}}(T)=C_{\psi_{x}^{*}Q}(T)$, where $U$ (resp, $Q$) is the universal subbundle (resp. quotient bundle) on $\mathbb{G}(l,2n-1)$ and $\psi_{x}$ is the morphism as in (3.2). So we have
$a_{l+1}X_{2}^{l+1}=(-1)^{l+1}c_{l+1}(\psi_{x}^{*}U)=c_{l+1}(\psi_{x}^{*}U^{\vee})~{}\text{and}~{}$
$b_{2n-1-l}X_{2}^{2n-1-l}=(-1)^{2n-1-l}c_{2n-1-l}(\psi_{x}^{*}Q)=c_{2n-1-l}(\psi_{x}^{*}Q)~{}(\text{since $l$ is odd,~{}$2n-1-l$ is even}).$
By [9, Proposition 2.1], $c_{l+1}(\psi_{x}^{*}U^{\vee})$ and $c_{2n-1-l}(\psi_{x}^{*}Q)$ are numerically non-negative, hence both $a_{l+1}$ and $b_{2n-1-l}$ are non-negative. As $-a$ is $a_{l+1}b_{2n-1-l}$ and $a$ is not 00, we have $-a>0$. So for any $t\in\mathbb{R}$, $t^{2n}-a$ is bigger than 00. In other words, $S_{1}(t,0,1)$ and $S_{2}(t,0,1)$ are non-zero for any $t\in\mathbb{R}$. By the equations $E(T,1)-aR(1,0)=S_{1}(T,1,0)S_{2}(T+1,1,0)=S_{1}(T,0,1)S_{2}(T+1,0,1)$, for any $t\in\mathbb{R}$, $E(t,1)-aR(1,0)$ is not 00. ∎
Proposition 3.9.
When $l$ is odd and $n$ is at least $3$, the approximate solution $\frac{1}{a}E(T,X_{1})$ or $\frac{1}{a}E(T-X_{1},X_{1})$ associated with (*3) is a primitive approximate solution.
Proof.
Recall the equation $\frac{1}{a}E(t,1)-R(1,z)=\frac{1}{a}S_{1}(t,1,z)S_{2}(t+1,1,z)$. Note that $l$ is odd implies $l+1\geq 2n-1-l>1$. So by Lemma 3.3, we have $\frac{1}{a}e_{2n}=0$ or $\frac{1}{a}e_{2n}=1$. And when $\frac{1}{a}e_{2n}$ is 00, $\frac{1}{a}E(0,1)-R(1,\exp\frac{2i\pi}{2n+2})$ vanishes; when $\frac{1}{a}e_{2n}$ is $1$, $\frac{1}{a}E(0,1)-R(1,\exp\frac{2i\pi}{2n})$ vanishes.
Suppose that $n$ is $3$. When $\frac{1}{a}e_{2n}$ is 00, we have $\frac{1}{a}S_{1}(0,1,z)S_{2}(1,1,z)=-(z^{6}+z^{4}+z^{2}+1)$. The prime factorization of $z^{6}+z^{4}+z^{2}+1$ over $\mathbb{Q}[z]$ is $(z^{4}+1)(z^{2}+1)$. Note $\deg(S_{1})\geq\deg(S_{2})$ and $\deg(S_{2})>1$, we must have $S_{1}(0,1,z)=\lambda(z^{4}+1)$ for some $\lambda\in\mathbb{Q}$. Then $\exp\frac{2i\pi}{8}$ is a root of $S_{1}(0,1,z)$ and hence $\frac{1}{a}E(T,X_{1})$ is primitive. When $\frac{1}{a}e_{2n}$ is $1$, we have $\frac{1}{a}S_{1}(0,1,z)S_{2}(1,1,z)=-(z^{6}+z^{4}+z^{2})=-z^{2}(z^{4}+z^{2}+1)=-z^{2}(z^{2}+z+1)(z^{2}-z+1)$. Note that $S_{1}$ is symmetrical in $X_{1},X_{2}$ and $\deg(S_{1})$ is at least $\deg(S_{2})$, we have $S_{1}(0,1,z)=\lambda(z^{4}+z^{2}+1)$. Then $\exp\frac{2i\pi}{6}$ is a root of $S_{1}(0,1,z)$ and hence $\frac{1}{a}E(T,X_{1})$ is also primitive.
Suppose now $n$ is at least $4$ and that $\frac{1}{a}E(T,X_{1})$ is not a primitive solution. Then we have
$S_{2}(1,1,\exp\frac{2i\pi}{2n+2})=0\text{ or }S_{2}(1,1,\exp\frac{2i\pi}{2n})=0.$ (3.3)
We now show that $\frac{1}{a}E(T-X_{1},X_{1})$ is a primitive solution. First we have the following inequalities:
$\frac{\pi}{2n}<\frac{2\pi}{2n+2}<\frac{2\pi}{2n}<\frac{3\pi}{2n}<\frac{4\pi}{2n+2}.$
(Note that for the last inequality, we use the condition $n\geq 4$). Since the roots of $S_{2}(0,1,z)$ satisfy $z^{2n}=1$ or $z^{2n+2}=1$, it suffices to show that $S_{2}(0,1,z)$ has a non-zero root $y_{0}$ with argument satisfying $\frac{\pi}{2n}<\arg y_{0}<\frac{3\pi}{2n}$. By the above inequalities, we have $y_{0}=\frac{2i\pi}{2n+2}$ or $y_{0}=\frac{2i\pi}{2n}$. Then $\frac{1}{a}E(T-X_{1},X_{1})$ is a primitive approximate solution.
It is enough to show that for any $t\in\mathbb{R}$, $S_{2}(t+1,1,z)$ as a polynomial of $z$ has a non-zero root with argument in $(\frac{\pi}{2n},\frac{3\pi}{2n})$. Denote by $K$ the set $\{t\in\mathbb{R}~{}|~{}\exists~{}r(t)(\neq 0)\in\mathbb{C},~{}S_{2}(t+1,1,r(t))=0\text{ with }\frac{\pi}{2n}<\arg r(t)<\frac{3\pi}{2n}\}$. By (3.3), we have $0\in K$. Note that $K$ is open by construction, if we can show $K$ is closed, then $K$ is $\mathbb{R}$. Let $t_{0}$ be a limit point of $K$. Then $S_{2}(t_{0}+1,1,z)$ has a root $r(t_{0})$ which is a limit point of $\{r(t)~{}|~{}t\in K\}$. By Lemma 3.8, for any $t\in\mathbb{R}$, as a polynomial in $z,$ the roots of $\frac{1}{a}E(t,1)-R(1,z)$ are non-zero. So the roots of $S_{2}(t+1,1,z)$ are also non-zero. In particular, $r(t_{0})$ is not zero. Suppose that $t_{0}\notin K$, then we have $\arg r(t_{0})=\frac{m\pi}{2n}$, where $m$ is $1$ or $3$. We may assume $r(t_{0})=\rho\exp\frac{im\pi}{2n}$, where $\rho$ is a positive real number. Then $\frac{1}{a}E(t_{0},1)-R(1,\rho\exp\frac{im\pi}{2n})$ is 00. To get a contradiction, we show for any $d\in\mathbb{R}$, we have $R(1,\rho\exp\frac{im\pi}{2n})+d\neq 0$. If $R(1,\rho\exp\frac{im\pi}{2n})+d=0$, we have the following identities:
$(R(1,\rho\exp\frac{im\pi}{2n})+d)(\rho^{2}\exp\frac{2im\pi}{2n}-1)$
$=$ $R(1,\rho\exp\frac{im\pi}{2n})(\rho^{2}\exp\frac{2im\pi}{2n}-1)+d(\rho^{2}\exp\frac{2im\pi}{2n}-1)$
$=$ $\rho^{2n+2}\exp\frac{im(2n+2)\pi}{2n}-1+d\rho^{2}\exp\frac{2im\pi}{2n}-d$
$=$ $(d\rho^{2}-\rho^{2n+2})\exp\frac{im\pi}{n}-(d+1)=0.$
Since $n$ is bigger than $3$ and $m$ is $1$ or $3$, $\exp\frac{im\pi}{n}$ is not real. We must have $d+1=0$ and $d\rho^{2}-\rho^{2n+2}=0$. But it implies $d\rho^{2}-\rho^{2n+2}=-\rho^{2}-\rho^{2n+2}=0$, which is absurd. ∎
Now we complete the proof for the case $l$ is odd. By Proposition 3.5 and Proposition 3.9, we have the following possibilities (note the coefficient of $T^{2n}$ in $E(T,X_{1})$ is $1$):
$E(T,X_{1})$ is $T^{2n}$ or $a\Sigma_{n}(\frac{T^{2}}{b_{1}},X_{1}^{2})$; $E(T-X_{1},X_{1})$ is $T^{2n}$ or $a\Sigma_{n}(\frac{T^{2}}{b_{2}},X_{1}^{2})$, where $b_{i}(\in\mathbb{Q})$ satisfy $b_{i}^{n}=a$$(i=1,2)$.
Proposition 3.10.
The above possibilities are all impossible.
Proof.
If $E(T,X_{1})$ is $T^{2n}$, $T^{2n}-aR(X_{1},X_{2})$ is $S_{1}(T,X_{1},X_{2})S_{2}(T+X_{1},X_{1},X_{2})$. However, we have
$X_{1}-\exp\frac{2i\pi}{2n+2}X_{2}\mid R(X_{1},X_{2}),~{}(X_{1}-\exp\frac{2i\pi}{2n+2}X_{2})^{2}\nmid R(X_{1},X_{2}),~{}X_{1}-\exp\frac{2i\pi}{2n+2}X_{2}\nmid 1.$
By Eisenstein’s criterion, $T^{2n}-aR(X_{1},X_{2})$ is irreducible as the polynomial in the variable $T$ with coefficients in $\mathbb{Q}[X_{1},X_{2}]$. This leads to a contradiction. Similar arguments can be applied to the case $E(T-X_{1},X_{1})$ is $T^{2n}$.
If $E(T,X_{1})$ is $b_{1}^{n}\Sigma_{n}(\frac{T^{2}}{b_{1}},X_{1}^{2})(=\Sigma_{n}(T^{2},b_{1}X_{1}^{2}))$, then we have the equalities $E(T,X_{1})-aR(X_{1},X_{2})=\Sigma_{n}(T^{2},b_{1}X_{1}^{2})-\Sigma_{n}(b_{1}X_{1}^{2},b_{1}X_{2}^{2})=(T^{2}-b_{1}X_{2}^{2})\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})$ (for the last equality, see [2, Section 7.2] for example). We will make use of the following claim, whose proof will be given in Proposition A.3.
Claim: $\Sigma_{n-1}(T^{2},X_{1}^{2},X_{2}^{2})$ is irreducible in $\mathbb{C}[T,X_{1},X_{2}]$.
As $b_{1}$ is not 00, $\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})$ is irreducible. We have $(T^{2}-b_{1}X_{2}^{2})\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})=S_{1}(T,X_{1},X_{2})S_{2}(T+X_{1},X_{1},X_{2})$. Since $S_{1}(T,X_{1},X_{2})$ is symmetrical in $X_{1},X_{2}$ and the coefficient of $T^{2n}$ of $S_{1}$ is $1$, then $S_{1}(T,X_{1},X_{2})$ is $\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})$ and hence $S_{2}(T+X_{1},X_{1},X_{2})$ is $T^{2}-b_{1}X_{2}^{2}$. So $S_{2}(T,X_{1},X_{2})$ is $(T-X_{1})^{2}-b_{1}X_{2}^{2},$ which is not symmetrical in $X_{1},X_{2}.$ We get a contradiction.
If $E(T-X_{1},X_{1})$ is $b_{2}^{n}\Sigma_{n}(\frac{T^{2}}{b_{2}},X_{1}^{2})(=\Sigma_{n}(T^{2},b_{2}X_{1}^{2}))$, we have $(T^{2}-b_{2}X_{2}^{2})\Sigma_{n-1}(T^{2},b_{2}X_{1}^{2},b_{2}X_{2}^{2})=S_{1}(T-X_{1},X_{1},X_{2})S_{2}(T,X_{1},X_{2})$. Similarly, $S_{2}(T,X_{1},X_{2})$ is $\Sigma_{n-1}(T^{2},b_{2}X_{1}^{2},b_{2}X_{2}^{2})$. But $\deg(S_{2})$ is at most $\deg(S_{1})$ and $n$ is at least $3$, it is impossible. ∎
Now we prove our main theorem.
Theorem 3.11.
Every uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}(n\geq 3)$ splits.
Proof.
We have shown that there is no unsplit uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}$ for $n\geq 3$. Now let $E^{\prime}$ be a uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+2}$. For every smooth hyperplane section $\mathbb{Q}^{2n+1}\hookrightarrow\mathbb{Q}^{2n+2}$, the restriction $E^{\prime}|_{\mathbb{Q}^{2n+1}}$ splits. Then we conclude by [6, Corollary 3.3]. ∎
Given a generalised Grassmannian $X$, we denote by $\mu(X)$ the maximum positive integer verifying that every uniform vector bundle whose rank is at most $\mu(X)$ splits. And we call $\mu(X)$ the splitting threshold for uniform vector bundles on $X$. For example, $\mu(\mathbb{P}^{n})$ is $n-1$. The main theorem in [3] tells us that for most generalised Grassmannians $X$, $\mu(X)$ is equal to $e.d.(\mathrm{VMRT})$ (see [3, Page 3, Table 2]). Note that for $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}$, their $e.d.(\mathrm{VMRT})$ are all $12n-$ 1.
Remark 3.12.
  1. (1)
    Theorem 3.11 reveals that the splitting thresholds for uniform vector bundles on $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}(n\geq 3)$ are at least $2n$. Actually, $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}(n\geq 3)$ are the first known examples such that $\mu(X)$ is bigger than $e.d.(\mathrm{VMRT})$.
  2. (2)
    The threshold $\mu(\mathbb{Q}^{2n+1})$ is $2n$, as the tangent bundle $T_{\mathbb{Q}^{2n+1}}$ is unsplit. Moreover, the splitting type of $T_{\mathbb{Q}^{2n+1}}$ is $(2,1,\dots,1,0)$.
In [3], the authors classify the unsplit uniform bundles of minimal ranks on the generalised Grassmannians $B_{n}/P_{k}~{}(2\leq k<\frac{2n}{3})$, $B_{n}/P_{n-2}$, $B_{n}/P_{n-1}$ and $D_{n}/P_{k}~{}(2\leq k<\frac{2n-2}{3})$, $D_{n}/P_{n-3}$, $D_{n}/P_{n-2}$. As direct corollaries of Theorem 3.11, we can give further classification results for uniform bundles on $B_{n}/P_{k}~{}(\frac{2n}{3}\leq k\leq n-3)$ and $D_{n}/P_{k}~{}(\frac{2n-2}{3}\leq k\leq n-4)$.
Corollary 3.13.
Let $X$ be $B_{n}/P_{k}$, where $k$ is $\frac{2n}{3}$ and $k$ is at least $6$. Let $E$ be a uniform vector bundle on $X$ of rank $r$.
  • If $r$ is smaller than $k$, then $E$ is a direct sum of line bundles.
  • If $r$ is $k$, then $E$ is either a direct sum of line bundles or $E_{\lambda_{1}}\otimes L$ or $E_{\lambda_{1}}^{\vee}\otimes L$ for some line bundle $L$, where $E_{\lambda_{1}}$ is the irreducible homogeneous bundle corresponding to the highest weight $\lambda_{1}$.
Proof.
Note that the $e.d.(\mathrm{VMRT})$ of $X$ is $k-1(=2n-2k-1)$. By [3, Theorem 1.1 (1)], the first assertion follows. For the case $r=k$, since $12n-2k+1=k+$ 1 is at least $7$, $E|_{\mathbb{Q}^{2n-2k+1}}$ splits by Theorem 3.11. Suppose $E$ is unsplit, then $E|_{\mathbb{P}^{k}}$ is also unsplit. Then the second assertion follows from [3, Proposition 4.4]. ∎
Corollary 3.14.
Let $X$ be $B_{n}/P_{k}$ with $\frac{2n}{3}<k\leq n-3$. Every uniform bundle of rank $2n-2k$ on $X$ splits.
Proof.
If $E$ is a uniform bundle of rank $2n-2k$ on $X$, then $E|_{\mathbb{P}^{k}}$ splits, as $2n-2k$ is smaller than $k$. On the other hand, since $k$ is at most $n-3$ and hence $12n-2k+$ 1 is at least $7$, $E|_{\mathbb{Q}^{2n-2k+1}}$ splits by Theorem 3.11. Because any $2$-plane in $X$ is contained in a $\mathbb{P}^{k}$ or $\mathbb{Q}^{2n-2k+1}$, $E$ splits by [1, Corollary 3.6]. ∎
Similar to the proofs of the above corollaries, we can prove the following results.
Corollary 3.15.
Let $X$ be $D_{n}/P_{k}$, where $k$ is $\frac{2n-2}{3}$ and $k$ is at least $6$. Let $E$ be a uniform vector bundle on $X$ of rank $r$.
  • If $r$ is smallet than $k$, then $E$ is a direct sum of line bundles.
  • If $r$ is $k$, then $E$ is either a direct sum of line bundles or $E_{\lambda_{1}}\otimes L$ or $E_{\lambda_{1}}^{\vee}\otimes L$ for some line bundle $L$, where $E_{\lambda_{1}}$ is the irreducible homogeneous bundle corresponding to the highest weight $\lambda_{1}$.
Corollary 3.16.
Let $X$ be $D_{n}/P_{k}$ with $\frac{2n-2}{3}<k\leq n-4$. Every uniform bundle of rank $22n-2k-$ 2 on $X$ splits.
Appendix A Classification of primitive approximate solutions
We use methods in [2] to classify primitive approximate solutions. Let $P(T,X_{1})=\sum_{k=0}^{2n}p_{k}X_{1}^{k}T^{2n-k}$ be a primitive approximate solution.
Proposition A.1.
If $p_{2n}$ is 00, then $P(T,X_{1})$ is $bT^{2n}$ for some rational number $b$.
Proof.
Let $S_{0}(T,X_{1},X_{2})$ be a symmetrical divisor of $P(T,X_{1})-R(X_{1},X_{2})$ such that for a $2(n+1)$-th primitive unit root $y_{0}$, we have $S_{0}(0,1,y_{0})=0$. Since $y_{0}$ is a simple root of $P(0,1)-R(1,z)$, $y_{0}$ is also a simple root of $S_{0}(0,1,z)$. Therefore, by the implicit function theorem, there is a germ of holomorphic function $y(x)$ at a neighborhood of $x=0$ satisfying
$S_{0}(x(1+y(x)),1,y(x))=0\text{ and }y(0)=y_{0}.$
We now show by induction that for $m=1,\dots,2n-1$, we have $y^{(m)}(0)=p_{2n-m}=0$. As $S_{0}$ is symmetrical in $X_{1}$ and $X_{2}$, we have
$P(x(1+y(x)),1)-R(y(x),1)=0,$ (1)
$P(x(1+y(x)),y(x))-R(y(x),1)=0.$ (2)
By taking the derivatives of (1) and (2) at $x=0$ and note that $p_{2n}$ is 00, we obtain
$p_{2n-1}(1+y_{0})-R^{\prime}(y_{0},1)y^{\prime}(0)=0,$ (1’)
$p_{2n-1}(1+y_{0})y_{0}^{2n-1}-R^{\prime}(y_{0},1)y^{\prime}(0)=0.$ (2’)
From $R(y,1)(y^{2}-1)=y^{2n+2}-1$, we have $R^{\prime}(y,1)(y^{2}-1)+R(y,1)\cdot(2y)=(2n+2)y^{2n+1}$. As $R(y_{0},1)$ is 00 and $y_{0}$ is not ±$\pm 1$, $R^{\prime}(y_{0},1)$ is $\frac{(2n+2)y_{0}^{2n+1}}{y_{0}^{2}-1}(\neq 0)$. From (1’) and (2’), we get $p_{2n-1}(1+y_{0})(y_{0}^{2n-1}-1)=0$. Since $y_{0}$ is primitive, $y_{0}^{2n-1}-1$ and $1+y_{0}$ are not zero. We have $p_{2n-1}=0$. By (1’) and note $R^{\prime}(y_{0},1)\neq 0$, $y^{\prime}(0)$ is 00.
For the case $m\geq 2$, we assume by induction that $y^{(m^{\prime})}(0)=p_{2n-m^{\prime}}=0$ for $m^{\prime}<m$. Then the $m$-th derivatives of (1) and (2) satisfy the following equations:
$m!p_{2n-m}(1+y_{0})^{m}-R^{\prime}(y_{0},1)y^{(m)}(0)=0,$ ($1^{m}$)
$m!p_{2n-m}(1+y_{0})^{m}y_{0}^{2n-m}-R^{\prime}(y_{0},1)y^{(m)}(0)=0.$ ($2^{m}$)
Since $y_{0}$ is primitive, $y_{0}^{2n-m}-1$ is not zero. We have $y^{(m)}(0)=p_{2n-m}=0$ as above. ∎
Proposition A.2.
If $p_{2n}$ is $1$, then $P(T,X_{1})$ is $\Sigma_{n}(bT^{2},X_{1}^{2})$ for some $b\in\mathbb{Q}$.
Proof.
Let $S_{+}(T,X_{1},X_{2})$ be a symmetrical divisor of $P(T,X_{1})-R(X_{1},X_{2})$ such that for a $2n$-th primitive unit root $y_{0}$, we have $S_{+}(0,1,y_{0})=0$. Note the equation $\Sigma_{n}(p_{2n-2}T^{2},X_{1}^{2})-R(X_{1},X_{2})=(p_{2n-2}T^{2}-X_{2}^{2})\Sigma_{n-1}(p_{2n-2}T^{2},X_{1}^{2},X_{2}^{2})$ (see [2, Section 7.2] for example). Let $S_{-}(T,X_{1},X_{2})$ be $\Sigma_{n-1}(p_{2n-2}T^{2},X_{1}^{2},X_{2}^{2})$, then $S_{-}(0,1,y_{0})$ is also 00. Denote $P(T,X_{1})$ by $P_{+}(T,X_{1})$ and denote $\Sigma_{n}(p_{2n-2}T^{2},X_{1}^{2})$ by $P_{-}(T,X_{1})$, we are going to show that $P_{+}$ equals $P_{-}$. Let $y_{\pm}(x)$ be germs of holomorphic functions satisfying
$S_{\pm}(x(1+y_{\pm}(x)),1,y_{\pm}(x))=0\text{ and }y_{\pm}(0)=y_{0}.$
As $S_{\pm}$ is symmetrical in $X_{1}$ and $X_{2}$, we have equations:
$P_{\pm}(x(1+y_{\pm}(x)),1)-R(y_{\pm}(x),1)=0,$ ($1_{\pm}$)
$P_{\pm}(x(1+y_{\pm}(x)),y_{\pm}(x))-R(y_{\pm}(x),1)=0.$ ($2_{\pm}$)
Let $p^{+}_{2n-m}$ (resp. $p^{-}_{2n-m}$) be the coefficient of $X_{1}^{2n-m}T^{m}$ in $P_{+}(T,X_{1})$ (resp. $P_{-}(T,X_{1})$). We prove $p^{+}_{2n-m}=p^{-}_{2n-m}$ and $y^{(m)}_{+}(0)=y^{(m)}_{-}(0)$ by induction for $m$ $(0\leq m\leq 2n)$.
When $m$ is 00, we have $p_{2n}^{+}=p_{2n}^{-}=1$, $y_{+}(0)=y_{-}(0)=y_{0}$. We next show $p_{2n-1}^{+}=p_{2n-1}^{-}=0$ and $y^{\prime}_{+}(0)=y^{\prime}_{-}(0)=0$. Taking the derivatives of ($1_{\pm}$) and ($2_{\pm}$) at $x=0$, we get
$p_{2n-1}^{\pm}(1+y_{0})-R^{\prime}(y_{0},1)y^{\prime}_{\pm}(0)=0,$ ($1^{\prime}_{\pm}$)
$p_{2n-1}^{\pm}(1+y_{0})y_{0}^{2n-1}+2ny_{0}^{2n-1}y^{\prime}_{\pm}(0)-R^{\prime}(y_{0},1)y^{\prime}_{\pm}(0)=0.$ ($2^{\prime}_{\pm}$)
From $R(y,1)(y^{2}-1)=y^{2n+2}-1$, we have $R^{\prime}(y,1)(y^{2}-1)+R(y,1)\cdot(2y)=(2n+2)y^{2n+1}$. Note $R(y_{0},1)=1$ (as $y_{0}$ is a $2n$-th primitive unit root). Substituting it in the above equation, we have
$2ny_{0}-y_{0}^{2}R^{\prime}(y_{0},1)=-R^{\prime}(y_{0},1).$ (†)
Multiplying ($2^{\prime}_{\pm}$) by $y_{0}^{2}$ and using the relation (A), one has
$p_{2n-1}^{\pm}(1+y_{0})y_{0}-R^{\prime}(y_{0},1)y_{\pm}^{\prime}(0)=0.$ ($y_{0}^{2}\cdot 2^{\prime}_{\pm}$)
$p_{2n-1}^{\pm}(1+y_{0})-R^{\prime}(y_{0},1)y_{\pm}^{\prime}(0)=0,$ ($1^{\prime}_{\pm}$)
Since $y_{0}$ is primitive, $\left|\begin{array}[]{ccc}y_{0}(1+y_{0})&-R^{\prime}(y_{0},1)\\ 1+y_{0}&-R^{\prime}(y_{0},1)\end{array}\right|=(1-y_{0})(1+y_{0})R^{\prime}(y_{0},1)$ is not 00. Then one obtains $p_{2n-1}^{\pm}=y_{\pm}^{\prime}(0)=0$.
When $m$ is at least $2$, by induction, we may assume $p^{+}_{2n-m^{\prime}}=p^{-}_{2n-m^{\prime}}$ and $y_{+}^{(m^{\prime})}(0)=y_{-}^{(m^{\prime})}(0)$ for $m^{\prime}<m$. By taking the $m$-th derivatives of ($1_{\pm}$) and ($2_{\pm}$) at $x=0$, we have
$m!p_{2n-m}^{\pm}(1+y_{0})^{m}-R^{\prime}(y_{0},1)y^{(m)}_{\pm}(0)+T^{1}_{\pm}=0,$ ($1^{m}_{\pm}$)
$m!p_{2n-m}^{\pm}(1+y_{0})^{m}y_{0}^{2n-m}+2ny_{0}^{2n-1}y^{(m)}_{\pm}(0)-R^{\prime}(y_{0},1)y^{(m)}_{\pm}(0)+T^{2}_{\pm}=0,$ ($2^{m}_{\pm}$)
where $T^{i}_{\pm}$ are the remaining terms satisfying $T^{1}_{-}=T^{1}_{+}$ and $T^{2}_{-}=T^{2}_{+}$. For the case $m=2$, by construction, we automatically have $p_{2n-2}^{-}=p_{2n-2}=p_{2n-2}^{+}$. Then one obtains $y_{+}^{(2)}(0)=y_{-}^{(2)}(0)$ from ($1^{m}_{\pm}$). Suppose now $m$ is at least $3$, multiplying ($2^{m}_{\pm}$) by $y_{0}^{2}$ and using (A), we have
$m!p_{2n-m}^{\pm}(1+y_{0})^{m}y_{0}^{2n-m+2}-R^{\prime}(y_{0},1)y_{\pm}^{(m)}(0)+y_{0}^{2}T^{2}_{\pm}=0.$ ($y_{0}^{2}\cdot 2^{m}_{\pm}$)
Note that $\left|\begin{array}[]{ccc}m!(1+y_{0})^{m}y_{0}^{2n-m+2}&-R^{\prime}(y_{0},1)\\ m!(1+y_{0})^{m}&-R^{\prime}(y_{0},1)\end{array}\right|=m!(1-y_{0}^{2n-m+2})(1+y_{0})^{m}R^{\prime}(y_{0},1)$ is not zero for $m\geq 3$. By solving the system of linear equations {($y_{0}^{2}\cdot 2^{m}_{\pm}$), ($1^{m}_{\pm}$)} (view $p_{2n-m}^{\pm}$ and $y_{\pm}^{(m)}(0)$ as indeterminate), we have $p_{2n-m}^{+}=p_{2n-m}^{-}$ and $y_{-}^{(m)}(0)=y_{+}^{(m)}(0)$. ∎
Proposition A.3.
$\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})$ is irreducible in $\mathbb{C}[T,X_{1},X_{2}]$.
Proof.
It suffices to show that the variety $V:=\{(T,X_{1},X_{2})|\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})=0\}\subset\mathbb{C}^{3}$ is smooth on $\mathbb{C}^{3}\backslash\{0\}$. By [2, Lemma 7.2], the variety defined by $\Sigma_{n}(T,X_{1},X_{2})=0$ is smooth on $\mathbb{C}^{3}\backslash 0$. Note that the map $(T,X_{1},X_{2})\mapsto(T^{2},X_{1}^{2},X_{2}^{2})$ is a local isomorphism outside the locus defined by $TX_{1}X_{2}=0$, $V$ is smooth on $\mathbb{C}^{3}\backslash\{TX_{1}X_{2}=0\}$.
Now suppose $(t,u,v)(\neq(0,0,0))$ is a singular point of $V$, then one of $t,u$ and $v$ is 00. By symmetry, we assume that $t$ is 00. Then there are equations
$\frac{\partial}{\partial T}\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})|_{(0,u,v)}=\frac{\partial}{\partial X_{1}}\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})|_{(0,u,v)}=\frac{\partial}{\partial X_{2}}\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})|_{(0,u,v)}=0.$
Furthermore, we have equations:
$\frac{\partial}{\partial X_{1}}\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})|_{(0,u,v)}=\frac{\partial}{\partial X_{1}}\Sigma_{n}(X_{1}^{2},X_{2}^{2})|_{(u,v)}=0,$
$\frac{\partial}{\partial X_{2}}\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})|_{(0,u,v)}=\frac{\partial}{\partial X_{2}}\Sigma_{n}(X_{1}^{2},X_{2}^{2})|_{(u,v)}=0.$
Without loss of generality, we may assume $v$ is not 00. As $\Sigma_{n}(u^{2},v^{2})(=\Sigma_{n}(0,u^{2},v^{2}))$ vanishes, $u/v$ is a root of $\Sigma_{n}(z^{2},1)=0$ with multiplicity greater than $2$. But $\Sigma_{n}(z^{2},1)=\frac{z^{2n+2}-1}{z^{2}-1}$ has no multiple roots, which is a contradiction. ∎