### v2

# uniform bundles on quadrics

Abstract.
We show that there exist only constant morphisms from $\mathbb{Q}^{2n+1}(n\geq 1)$ to $\mathbb{G}(l,2n+1)$ if $l$ is even $(0<l<2n)$ and $(l,2n+1)$ is not $(2,5)$. As an application, we prove on $\mathbb{Q}^{2m+1}$ and $\mathbb{Q}^{2m+2}(m\geq 3)$, any uniform bundle of rank at most $2m$ splits, which improves the upper bound of splitting for uniform bundles obtained by Kachi and Sato [7]. We classify all unsplit uniform bundles of minimal rank on $B_{n}/P_{k}$ $(k=\frac{2n}{3},k\geq 6)$ and $D_{n}/P_{k}$ $(k=\frac{2n-2}{3},k\geq 6)$. We partially answer a conjecture of Ellia, which predicts that some uniform bundles of special splitting types on $\mathbb{P}^{n}$ necessarily split and we find some restrictions on the splitting types of unsplit uniform bundles of minimal rank.
1.Xinyi Fang is supported by National Natural Science Foundation of China (Grant No. 12471040)
2.Duo Li is supported by National Natural Science Foundation of China (Grant No. 12001547)
Keywords: uniform bundle; quadric; generalized Grassmannian.
MSC: 14M15; 14M17; 14J60.
1. Introduction
In this article, we assume all the varieties are defined over $\mathbb{C}$. We assume that $X$ is a rational homogeneous variety of Picard number $1$ and we call $X$ a generalized Grassmannian for short. It is well known that $X$ is swept by lines. Let $E$ be a vector bundle on $X$. We consider its restriction $E|_{L}(\simeq\mathcal{O}_{L}(a_{1}(L)\oplus\cdots\oplus\mathcal{O}_{L}(a_{r}(L)))$ to any line $L\subseteq X$. If the splitting type $(a_{1}(L),\dots,a_{r}(L))$ of $E|_{L}$ is independent of the choice of $L$, $E$ is called a uniform bundle. We say a vector bundle splits if it can be decomposed as a direct sum of line bundles. We say a vector bundle does not split or a vector bundle is unsplit if it cannot be decomposed as a direct sum of line bundles.
In [5], Grothendieck shows that every vector bundle on a projective line splits. However, for higher dimensional projective spaces, the splitting of vector bundles is far more intricate. In [15], Van de Ven studies the splitting property of uniform bundles. He demonstrates that every uniform bundle of rank $2$ on $\mathbb{P}^{n}$ $(n\geq 3)$ splits and every uniform bundle of rank $2$ on $\mathbb{P}^{2}$ is isomorphic to $\mathcal{O}_{\mathbb{P}^{2}}(a)\oplus\mathcal{O}_{\mathbb{P}^{2}}(b)$ or $T_{\mathbb{P}^{2}}(a)$ for some integers $a$ and $b$. The subsequent work by Sato [12] and Elencwajg, Hirschowitz, Schneider [2] extends these results: every uniform bundle on $\mathbb{P}^{n}$ of rank smaller than $n$ splits, while every unsplit uniform bundle on $\mathbb{P}^{n}$ of rank $n$ is isomorphic to $T_{\mathbb{P}^{n}}(a)$ or $\Omega_{\mathbb{P}^{n}}(b)$ for some integers $a$ and $b$.
Motivated by these advances, we address two central problems for uniform bundles on a generalized Grassmannian $X$.
  • Problem $1$: Determine the splitting threshold $\mu(X)$ such that any uniform bundle of rank at most $\mu(X)$ splits and there exists an unsplit uniform vector bundle of rank $\mu(X)+1$.
  • Problem $2$: Classify all uniform bundles of rank $\mu(X)+1$.
Let $\mathbb{G}(k-1,n-1)$ be the Grassmannian of $k$-dimensional subspaces of $\mathbb{C}^{n}$. In [6], Guyot solves the above problems for Grassmannians $X=\mathbb{G}(k-1,n-1)(k\leq n-k)$. For orthogonal Grassmannians $OG(n,2n+1)$, Muñoz, Occhetta and Solá Conde [8] achieve similar results.
Let $\mathbb{Q}^{n}$ be a projective smooth quadric of dimension $n$, with $n\geq 5$ odd (resp. even). Kachi and Sato [7, Theorem 4.1] prove that any uniform vector bundle on $\mathbb{Q}^{n}$ of rank at most $n-2$ (resp. $n-3$) splits. Muñoz, Occhetta and Solá Conde provide an alternative proof via a general splitting criterion for low-rank uniform bundles on varieties covered by lines (see [8, Corollary 3.3]). However we will demonstrate their upper bounds are not optimal. Actually, we show that for $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}$ $(n\geq 3)$, every uniform bundle of rank $2n$ splits.
In our recent work [4], we systematically address Problem $1$ and Problem $2$ for all generalized Grassmannians. A key insight is establishing a connection between the geometry of VMRT (variety of minimal rational tangents) and the splitting behavior of uniform bundles. Specifically, we demonstrate that for an arbitrary generalized Grassmannian $X$, a uniform bundle $E$ of rank at most the so-called $e.d.(\mathrm{VMRT})$ (for the definition of $e.d.(\mathrm{VMRT})$, see [4, Section $3$]) must split. Furthermore, for the majority of generalized Grassmannians, the bound $e.d.(\mathrm{VMRT})$ coincides with $\mu(X)$. Our approach diverges from Guyot’s algebraic strategy by adopting a more geometric and direct methodology. Notably, for most generalized Grassmannians whose VMRTs are products of several irreducible varieties, we reduce Problem $2$ to the classification of uniform bundles on projective spaces and quadrics, see [4, Proposition $4.4$]. This reduction highlights the pivotal role of quadrics in resolving the problem.
Given a generalized Grassmannian $X$, assume $\mu(X)$ is $e.d.(\text{VMRT})$. Let $E$ be an unsplit uniform bundle of minimal rank on $X$, that is, the rank of $E$ is $\mu(X)+1$. In [4, Proposition 3.8], we show that, up to twisting by a suitable line bundle, the splitting type of $E$ is $(1,\dots,1,0,\dots,0)$. Inspired by this result, when $\mu(X)$ is not necessarily $e.d.(\text{VMRT})$, we show that there are some restrictions on the splitting type of $E$. To be concrete, if the splitting type of $E$ is $(a_{1},a_{2},\dots,a_{r})$ $(a_{1}\geq a_{2}\geq\cdots\geq a_{r})$, then $max\{a_{i}-a_{i+1}|1\leq i\leq r-1\}$ is $1$. By using the same strategy, we give an affirmative partial answer to a conjecture of Ellia (see [3, Page 29, Conjecture]), which predicts that every uniform bundle on $\mathbb{P}^{n}$ of splitting type $(\underbrace{a_{1},\dots,a_{1}}_{l_{1}},\underbrace{a_{2},\dots,a_{2}}_{l_{2}},\dots,\underbrace{a_{k},\dots,a_{k}}_{l_{k}})$ with $a_{i}>a_{i+1}$ and $l_{i}\leq n-1$ for any $1\leq i\leq k$ necessarily splits. Motivated by Ellia’s conjecture, we propose a similar conjecture for any generalized Grassmannian as follows:
Conjecture: For an arbitrary generalized Grassmannian $X$, there exists a maximal positive integer $\nu(X)$ such that every uniform bundle of splitting type
$(\underbrace{a_{1},\dots,a_{1}}_{l_{1}},\underbrace{a_{2},\dots,a_{2}}_{l_{2}},\dots,\underbrace{a_{k},\dots,a_{k}}_{l_{k}})$
with $a_{i}>a_{i+1}$ and $l_{i}\leq\nu(X)$ for any $1\leq i\leq k$ necessarily splits.
In this article, we establish:
Main Results:
  1. (1)
    For $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}$ $(n\geq 3)$, every uniform bundle of rank $2n$ splits, with $\mu(\mathbb{Q}^{2n+1})=2n$.
  2. (2)
    For $B_{n}/P_{k}$ $(k=\frac{2n}{3})$ and $D_{n}/P_{k}$$(k=\frac{2n-2}{3})$, we classify all unsplit uniform bundles of minimal rank.
  3. (3)
    Give an affirmative partial answer to a conjecture of Ellia in a more general setting and find some restrictions on the splitting types of unsplit uniform bundles of minimal rank.
Note that for $\mathbb{Q}^{2n+1}$, $e.d.(\mathrm{VMRT})$ is $12n-$ 1 and the number $\mu(\mathbb{Q}^{2n+1})$ is $2n$. So this is the first known example such that the optimal upper bound $\mu(X)$ is bigger than $e.d.(\mathrm{VMRT})$.
This article is structured as follows:
  1. (1)
    Section $2$: Analyze morphisms from $\mathbb{Q}^{2n+1}$ to $\mathbb{G}(l,2n+1)$ and the main result of this section is Proposition 2.1.
  2. (2)
    Theorem 3.11: Prove splitting theorems for quadrics via relative Harder-Narasimhan filtrations and approximate solutions.
  3. (3)
    Corollaries 3.133.163.133.163.13–3.163.13 – 3.16: Extend classifications to generalized Grassmannians $B_{n}/P_{k}$ and $D_{n}/P_{k}$.
  4. (4)
    Section $4$: Prove some uniform bundles of special splitting types necessarily split.
2. Morphisms from quadrics to Grassmannians
Proposition 2.1.
There exist only constant morphisms from $\mathbb{Q}^{2n+1}(n\geq 1)$ to $\mathbb{G}(l,2n+1)$ if $l$ is even $(0<l<2n)$ and $(l,2n+1)$ is not $(2,5)$.
Proof.
We follow the proof of the main theorem in [14]. Let $H$ denote the cohomology class of a hyperplane on $\mathbb{Q}^{2n+1}$. The cohomology ring of $\mathbb{Q}^{2n+1}$ is
$H^{\bullet}(\mathbb{Q}^{2n+1},\mathbb{Z})=\mathbb{Z}\oplus\mathbb{Z}H\oplus\cdots\oplus\mathbb{Z}H^{n-1}\oplus\mathbb{Z}\frac{H^{n}}{2}\oplus\cdots\oplus\mathbb{Z}\frac{H^{2n+1}}{2}.$
Let $U$ (resp. $Q$) be the universal subbundle (resp. quotient bundle) on $\mathbb{G}(l,2n+1)$. Suppose that $f:\mathbb{Q}^{2n+1}\rightarrow\mathbb{G}(l,2n+1)$ is a non-constant morphism. Let $c_{i}$ and $d_{j}$ be rational numbers satisfying
$c_{i}(f^{*}U^{\vee})=c_{i}H^{i}\text{ for }1\leq i\leq l+1~{}\text{ and }~{}c_{j}(f^{*}Q)=d_{j}H^{j}\text{ for }1\leq j\leq 2n+1-l.$
By [13, Proposition 2.1 (i)], all coefficients $c_{i}(1\leq i\leq l+1)$ and $d_{j}(1\leq j\leq 2n+1-l)$ are non-negative. We note that for any $1\leq i,j<n$, $c_{i}$ and $d_{j}$ are integers. For any $i,j\geq n$, $2c_{i}$ and $2d_{j}$ are integers. Then from the exact sequence $0\rightarrow f^{*}U\rightarrow\mathcal{O}_{\mathbb{Q}^{2n+1}}^{\oplus 2n+2}\rightarrow f^{*}Q\rightarrow 0$, we get the equality of polynomials:
$(1-c_{1}t+c_{2}t^{2}+\cdots+(-1)^{l+1}c_{l+1}t^{l+1})(1+d_{1}t+\cdots+d_{2n+1-l}t^{2n+1-l})$ (2.1)
$=$ $1+(-1)^{l+1}c_{l+1}d_{2n+1-l}t^{2n+2}.$
If $c_{l+1}d_{2n+1-l}$ is 00, then from (2.1), we obtain both $c_{1}$ and $d_{1}$ are zero by induction, which implies that $f$ is constant. So we may assume the numbers $c_{1},d_{1},c_{l+1}$ and $d_{2n+1-l}$ are non-zero. By [13, Proposition 2.1 (ii)], all the rational numbers $c_{1},c_{2},\dots,c_{l+1}$ and $d_{1},d_{2},\dots,d_{2n+1-l}$ are positive.
Let $a$ be $2n+2]{c_{l+1}d_{2n+1-l}}$. We set $C_{i}:=\frac{c_{i}}{a^{i}}(1\leq i\leq l+1)$ and $D_{j}:=\frac{d_{j}}{a^{j}}(1\leq j\leq 2n+1-l)$. We note that if $a,C_{i}$ and $D_{j}$ are all positive integers, then by the same proof of [14, Page 204, Case (ii)], we can get a contradiction (for details, see Appendix A, Claim). Then it suffices to show that $a,C_{i}$ and $D_{j}$ are integers.
We first show that $a$ is an integer. Since $l$ is even, similar to the proof in [13, Lemma 3.3 (ii) and Case 1 of (iii)], we can show that $a$ is $\frac{c_{m+1}}{c_{m}}$ (see Appendix A, Corollary A.2), where $m$ is $\frac{l}{2}$. So $a$ is rational, we may assume $a=s/t$, where $s$ and $t$ are coprime positive integers. By definition, we have $a^{2n+2}=c_{l+1}d_{2n+1-l}$. Since $2c_{l+1}$ and $2d_{2n+1-l}$ are integers, $4a^{2n+2}$ is an integer. So $t^{2n+2}$ divides $4$, which implies that $t$ is $1$, as $n$ is at least $1$. Hence $a$ is an integer.
We now show that $C_{i}(=\frac{c_{i}}{a^{i}})$ and $D_{j}(=\frac{d_{j}}{a^{j}})$ are integers. Let $F_{1}(x)\cdots F_{k}(x)$ be the irreducible factorization of $1-x^{2n+2}$ over $\mathbb{Z}[x]$ with $F_{l}(0)=1(1\leq l\leq k)$. Then $F_{l}(x)$ is also irreducible over $\mathbb{Q}[x]$ by Gauss’s Lemma. Then
$(1-c_{1}t+c_{2}t^{2}+\cdots+(-1)^{l+1}c_{l+1}t^{l+1})(1+d_{1}t+\cdots+d_{2n+1-l}t^{2n+1-l})$
$=$ $1+(-1)^{l+1}c_{l+1}d_{2n+1-l}t^{2n+2}=1-a^{2n+2}t^{2n+2}=F_{1}(at)\cdots F_{k}(at).$
Note that $F_{i}(at)$ is also irreducible over $\mathbb{Q}[t]$. We have
$1-c_{1}t+c_{2}t^{2}+\cdots+(-1)^{l+1}c_{l+1}t^{l+1}=F_{i_{1}}(at)\cdots F_{i_{k_{1}}}(at)~{}\text{and}~{}$
$1+d_{1}t+\cdots+d_{2n+1-l}t^{2n+1-l}=F_{j_{1}}(at)\cdots F_{j_{k_{2}}}(at).$
Since the coefficients of $F_{l}(x)$ are integers, $C_{i}(=\frac{c_{i}}{a^{i}})$ and $D_{j}(=\frac{d_{j}}{a^{j}})$ are integers. ∎
3. Uniform bundles on $\mathbb{Q}^{2n+1}(n\geq 3)$
We are going to use the method in [2] to show that any uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}(n\geq 3)$ splits. We first fix some notations.
Let $E$ be a uniform bundle on $X(=\mathbb{Q}^{2n+1})(n\geq 3)$ of rank $2n$. Assume that $E$ does not split, by [4, Proposition $3.7$], we may assume that the splitting type of $E$ is
$(\underbrace{0,\dots,0}_{l+1},\underbrace{-1,\dots,-1}_{2n-l-1})\quad(l+1\geq 2n-l-1\geq 1).$
We denote the moduli of lines on $X$ and the corresponding universal family by:
${\mathcal{U}(=B_{n+1}/P_{1,2})}$${\mathcal{M}(=B_{n+1}/P_{2})}$${X(=B_{n+1}/P_{1}).}$$\scriptstyle{q}$$\scriptstyle{p}$
where $\mathcal{M}$ is the moduli of lines and $\mathcal{U}$ is the universal family. The relative Harder-Narasimhan (H-N) filtration of $p^{*}E$ induces an exact sequence:
$0\rightarrow E_{1}(=q^{*}G_{1})\rightarrow p^{*}E\rightarrow E_{2}(=q^{*}G_{2}\otimes p^{*}\operatorname{\mathcal{O}}_{X}(-1))\rightarrow 0,$ (3.1)
where $G_{1}$ (resp. $G_{2}$) is a vector bundle on $\mathcal{M}$ of rank $l+1$ (resp. $12n-l-$ 1). For each $x\in X$, the restriction of relative H-N filtration to $p^{-1}(x)$ induces a morphism
$\psi_{x}:p^{-1}(x)(\cong\mathbb{Q}^{2n-1})\rightarrow Gr(l+1,2n)(\cong\mathbb{G}(l,2n-1)).$ (3.2)
By [4, Lemma $2.1$], we have the following description of cohomology rings:
$H^{\bullet}(X,\mathbb{Q})=\mathbb{Q}[X_{1}]/(X_{1}^{2n+2}),$
$H^{\bullet}(\mathcal{U},\mathbb{Q})=\mathbb{Q}[X_{1},X_{2}]/(\Sigma_{i}(X_{1}^{2},X_{2}^{2})_{n\leq i\leq n+1}),$
$H^{\bullet}(\mathcal{M},\mathbb{Q})=\mathbb{Q}[X_{1}+X_{2},X_{1}X_{2}]/(\Sigma_{i}(X_{1}^{2},X_{2}^{2})_{n\leq i\leq n+1}).$
For a bundle $F$ of rank $r$, the Chern polynomial of $F$ is defined as
$C_{F}(T):=T^{r}-c_{1}(F)T^{r-1}+\cdots+(-1)^{r}c_{r}(F).$
Let $E(T,X_{1})=\sum_{k=0}^{2n}e_{k}X_{1}^{k}T^{2n-k}(\in\mathbb{Q}[X_{1},T])$ and $S_{i}(T,X_{1},X_{2})(\in\mathbb{Q}[X_{1}+X_{2},X_{1}X_{2},T])(i=1,2)$ be homogeneous polynomials representing $C_{p^{*}E}(T)$ and $C_{q^{*}G_{i}}(T)$ in the cohomology rings respectively. There are equations
$E(T,X_{1})=C_{p^{*}E}(T)~{}\text{and}~{}S_{i}(T,X_{1},X_{2})=C_{q^{*}G_{i}}(T)(i=1,2).$
Let $R(X_{1},X_{2})$ be the polynomial $\Sigma_{n}(X_{1}^{2},X_{2}^{2})$. By (3.1), we have an equation of Chern polynomials:
$E(T,X_{1})-aR(X_{1},X_{2})=S_{1}(T,X_{1},X_{2})S_{2}(T+X_{1},X_{1},X_{2}).$ (*)
If $a$ is 00, then both $S_{1}(T,X_{1},X_{2})$ and $S_{2}(T+X_{1},X_{1},X_{2})$ are polynomials only in variables $T$ and $X_{1}$. Since $S_{i}(T,X_{1},X_{2})(i=1,2)$ are symmetric in $X_{1}$ and $X_{2}$, we must have $S_{1}(T,X_{1},X_{2})=T^{l+1}$ and $S_{2}(T,X_{1},X_{2})=T^{2n-l-1}$. So $c_{1}(E_{1})$ and $c_{1}(E_{2})$ are 00, and $\psi_{x}$ is constant for each $x\in X$, which implies that $E$ splits (see, for example, [4, Proposition $3.5$]). Therefore, we have $a\neq 0$.
3.1. Approximate solutions
To solve the equation (*3), we use the concept of approximate solutions introduced in [2, Section 5]. The following definitions and propositions are basically from [2, Section 5] and the proofs are similar.
Definition 3.1.
A non-zero homogeneous polynomial $P(T,X_{1})$ with rational coefficients in variables $T$ and $X_{1}$ of degree $2n$ is called an approximate solution if $P(T,X_{1})-R(X_{1},X_{2})$ has a proper divisor $S(T,X_{1},X_{2})$ which is symmetric in $X_{1}$ and $X_{2}$. We call such a divisor a symmetric divisor.
Example 3.2.
By the equation (*3), both $\frac{1}{a}E(T,X_{1})$ and $\frac{1}{a}E(T-X_{1},X_{1})$ are approximate solutions. We call them approximate solutions associated with (*3).
In the following lemma, there are some restrictions on the coefficient of $X_{1}^{2n}$ for an arbitrary approximate solution $P(T,X_{1})$.
Lemma 3.3.
Let $P(T,X_{1})=\sum_{k=0}^{2n}p_{k}X_{1}^{k}T^{2n-k}$ be an approximate solution. Then one of the following holds.
$(1)$ Any symmetric divisor of $P(T,X_{1})-R(X_{1},X_{2})$ is of degree one.
$(2)$ The coefficient $p_{2n}$ is 00 and the zero set of $P(0,1)-R(1,z)$ is ($(\{z|z^{2n+2}-1=0\})\backslash\{1,-1\}$.
$(3)$ The coefficient $p_{2n}$ is $1$ and the zero set of $P(0,1)-R(1,z)$ is ($(\{z|z^{2n}-1=0\}\cup\{0\})\backslash\{1,-1\}$.
Proof.
Let $S(T,X_{1},X_{2})$ be a symmetric divisor of $P(T,X_{1})-R(X_{1},X_{2})$. Then $S(0,X_{1},X_{2})$ divides $p_{2n}X_{1}^{2n}-R(X_{1},X_{2})$. As $S(T,X_{1},X_{2})$ is symmetric in $X_{1}$ and $X_{2}$, $S(0,X_{1},X_{2})$ divides $p_{2n}X_{2}^{2n}-R(X_{1},X_{2}).$ Therefore $S(0,X_{1},X_{2})$ divides $p_{2n}(X_{1}^{2n}-X_{2}^{2n})$. By the equation $R(X_{1},X_{2})(X_{1}^{2}-X_{2}^{2})=X_{1}^{2n+2}-X_{2}^{2n+2}$, we have
$(X_{1}^{2}-X_{2}^{2})(p_{2n}X_{1}^{2n}-R(X_{1},X_{2}))+X_{2}^{2}(X_{1}^{2n}-X_{2}^{2n})=(p_{2n}-1)X_{1}^{2n}(X_{1}^{2}-X_{2}^{2}).$
If $p_{2n}$ is neither 00 nor $1$, we have $S(0,X_{1},X_{2})\mid X_{1}^{2n}-X_{2}^{2n}$ and $S(0,X_{1},X_{2})\mid X_{1}^{2n}(X_{1}-X_{2})(X_{1}+X_{2})$. Since $S(0,X_{1},X_{2})$ is symmetric in $X_{1}$ and $X_{2}$, $S(0,X_{1},X_{2})$ is $c(X_{1}+X_{2})$ for some $c\in\mathbb{Q}$. In particular, $\deg(S)$ is $1$.
If $p_{2n}$ is 00, then $P(0,1)-R(1,z)$ is $-\frac{z^{2n+2}-1}{z^{2}-1}(=-(z^{2n}+z^{2n-2}+\cdots+1))$. If $p_{2n}$ is $1$, then $P(0,1)-R(1,z)$ is $-z^{2}\frac{z^{2n}-1}{z^{2}-1}=(-(z^{2n}+z^{2n-2}+\cdots+z^{2}))$. Then the assertions follow immediately. ∎
Definition 3.4.
We call an approximate solution $P(T,X_{1})=\sum_{k=0}^{2n}p_{k}X_{1}^{k}T^{2n-k}$ a primitive approximate solution if $p_{2n}\in\{0,1\}$ and $P(T,X_{1})-R(X_{1},X_{2})$ has a symmetric divisor $S_{0}(T,X_{1},X_{2})$ such that there is a $2(n-p_{2n}+1)$-th primitive unit root $y_{0}$ satisfying $S_{0}(0,1,y_{0})=0$.
As in [2], we have the following classifications of primitive approximate solutions.
Proposition 3.5.
Let $P(T,X_{1})$ be a primitive approximate solutions. If $p_{2n}$ is 00, we have $P(T,X_{1})=bT^{2n}$ for some $b\in\mathbb{Q}$. If $p_{2n}$ is $1$, we have $P(T,X_{1})=\Sigma_{n}(bT^{2},X_{1}^{2})$ for some $b\in\mathbb{Q}$.
The proof of Proposition 3.5 is similar to that of [2, Proposition 6.1 and Proposition 6.2], we leave them in the Appendix (see Propositions B.1 and B.2).
3.2. The case $l$ is even
We first show that $l$ is not even and we will prove it by contradiction. Now suppose that $l$ is even. If $l$ is smaller than $22n-$ 2 and $(l,2n-1)$ is not $(2,5)$, the morphism $\psi_{x}$ (for the definition of $\psi_{x}$, see (3.2)) is constant for each $x\in X$ according to Proposition 2.1. Then $E$ splits. We exclude the remaining cases $l=2n-2$ and $(l,2n-1)=(2,5)$ by calculations.
Proposition 3.6.
There does not exist an unsplit uniform bundle of rank $2n$ whose splitting type is $(0,\dots,0,-1)$ on $\mathbb{Q}^{2n+1}(n\geq 3)$.
Proof.
Suppose $E$ is unsplit, by the same calculation as in [4, Theorem 4.3, Case I], we have $c_{1}(E_{2})=(X_{1}+X_{2})-X_{1}=X_{2}$. Let $f(X_{1},X_{2})$ be a homogeneous polynomial of degree $12n-$ 1 which is symmetric in $X_{1}$ and $X_{2}$ and represents $c_{2n-1}(E_{1})$. By comparing the coefficients of $T^{0}$ on the left and right sides of the equation (*3), we get
$f(X_{1},X_{2})X_{2}=e_{2n}X_{1}^{2n}-a(X_{1}^{2n}+X_{1}^{2n-2}X_{2}^{2}+\cdots+X_{2}^{2n}).$
Then we must have $e_{2n}=a$ and $f(X_{1},X_{2})=-a(X_{1}^{2n-2}X_{2}+X_{1}^{2n-4}X_{2}^{3}+\cdots+X_{2}^{2n-1})$, contradicting the assumption that $f$ is symmetric in $X_{1}$ and $X_{2}$. ∎
We now exclude the case $(l,2n-1)=(2,5)$.
Proposition 3.7.
There does not exist an unsplit uniform of rank $6$ whose splitting type is $(0,0,0,-1,-1,-1)$ on $\mathbb{Q}^{7}$.
Proof.
Suppose $E$ is unsplit. Then $a$ in the equation (*3) is not 00 and $\frac{1}{a}E(T,X_{1})=\frac{1}{a}\sum_{k=0}^{6}e_{k}X_{1}^{k}T^{6-k}$ is an approximate solution which has a symmetric divisor of degree $3$. By Lemma 3.3, we have $\frac{1}{a}e_{6}=0$ or $\frac{1}{a}e_{6}=1$. Let $f(X_{1},X_{2})$ be a homogeneous polynomial representing $c_{3}(E_{1})$.
If $\frac{1}{a}e_{6}$ is 00, we have $f(X_{1},X_{2})\mid R(X_{1},X_{2})(=X_{1}^{6}+X_{1}^{4}X_{2}^{2}+X_{1}^{2}X_{2}^{4}+X_{2}^{6})$. Since the prime factorization of $X_{1}^{6}+X_{1}^{4}X_{2}^{2}+X_{1}^{2}X_{2}^{4}+X_{2}^{6}$ over $\mathbb{Q}[X_{1},X_{2}]$ is $(X_{1}^{4}+X_{2}^{4})(X_{1}^{2}+X_{2}^{2})$, $R(X_{1},X_{2})$ has no divisor symmetric in $X_{1}$ and $X_{2}$ of degree $3$.
If $\frac{1}{a}e_{6}$ is $1$, then $f(X_{1},X_{2})$ divides $X_{1}^{6}-R(X_{1},X_{2})(=-X_{2}^{2}(X_{1}^{4}+X_{1}^{2}X_{2}^{2}+X_{2}^{4}))$. The prime factorization of $X_{1}^{4}+X_{1}^{2}X_{2}^{2}+X_{2}^{4}$ over $\mathbb{Q}[X_{1},X_{2}]$ is $(X_{1}^{2}+X_{1}X_{2}+X_{2}^{2})(X_{1}^{2}-X_{1}X_{2}+X_{2}^{2})$. Therefore, $X_{1}^{6}-R(X_{1},X_{2})$ has no divisor symmetric in $X_{1}$ and $X_{2}$ of degree $3$.
In both cases, we get contradictions. ∎
3.3. The case $l$ is odd
Suppose that $l$ is odd. We begin with a lemma.
Lemma 3.8.
When $l$ is odd, the equation $E(t,1)-aR(1,0)=0$ has no roots in $\mathbb{R}$.
Proof.
In the equation (*3): $E(T,X_{1})-aR(X_{1},X_{2})=S_{1}(T,X_{1},X_{2})S_{2}(T+X_{1},X_{1},X_{2})$, we let $X_{1}$ be 00 and let $X_{2}$ be $1$. Then we get an equation $T^{2n}-a=S_{1}(T,0,1)S_{2}(T,0,1)$. We write $S_{1}(T,0,X_{2})$ and $S_{2}(T,0,X_{2})$ as follows: $S_{1}(T,0,X_{2})=\sum_{i=0}^{l+1}a_{i}X_{2}^{i}T^{l+1-i}$ and $S_{2}(T,0,X_{2})=\sum_{j=0}^{2n-1-l}b_{j}X_{2}^{j}T^{2n-1-l-j}$, where $a_{i}(0\leq i\leq l+1)$ and $b_{j}(0\leq j\leq 2n-1-l)$ are rational numbers. Then $-a$ is $a_{l+1}b_{2n-1-l}$. We now wish to show $-a>0$. To this end, for any $x$ in $X$, we consider the embedding $i_{x}:p^{-1}(x)(=\mathbb{Q}^{2n-1})\hookrightarrow\mathcal{U}(=B_{n+1}/P_{1,2})$, which induces a morphism:
$i_{x}^{*}:H^{\bullet}(\mathcal{U},\mathbb{Q})\cong\mathbb{Q}[X_{1},X_{2}]/(\Sigma_{n}(X_{1}^{2},X_{2}^{2}),\Sigma_{n+1}(X_{1}^{2},X_{2}^{2}))\rightarrow\mathbb{Q}[X_{2}]/X_{2}^{2n}\cong H^{\bullet}(\mathbb{Q}^{2n-1},\mathbb{Q}).$
Under the above identifications, we have $S_{1}(T,0,X_{2})=C_{E_{1}|_{p^{-1}(x)}}(T)=C_{\psi_{x}^{*}U}(T)$ and $S_{2}(T,0,X_{2})=C_{E_{2}|_{p^{-1}(x)}}(T)=C_{\psi_{x}^{*}Q}(T)$, where $U$ (resp, $Q$) is the universal subbundle (resp. quotient bundle) on $\mathbb{G}(l,2n-1)$ and $\psi_{x}$ is the morphism as in (3.2). So we have
$a_{l+1}X_{2}^{l+1}=(-1)^{l+1}c_{l+1}(\psi_{x}^{*}U)=c_{l+1}(\psi_{x}^{*}U^{\vee})~{}\text{and}~{}$
$b_{2n-1-l}X_{2}^{2n-1-l}=(-1)^{2n-1-l}c_{2n-1-l}(\psi_{x}^{*}Q)=c_{2n-1-l}(\psi_{x}^{*}Q)~{}(\text{since $l$ is odd,~{}$2n-1-l$ is even}).$
By [13, Proposition 2.1], $c_{l+1}(\psi_{x}^{*}U^{\vee})$ and $c_{2n-1-l}(\psi_{x}^{*}Q)$ are numerically non-negative, hence both $a_{l+1}$ and $b_{2n-1-l}$ are non-negative. As $-a$ is $a_{l+1}b_{2n-1-l}$ and $a$ is not 00, we have $-a>0$. So for any $t\in\mathbb{R}$, $t^{2n}-a$ is bigger than 00. In other words, $S_{1}(t,0,1)$ and $S_{2}(t,0,1)$ are non-zero for any $t\in\mathbb{R}$. By the equations $E(T,1)-aR(1,0)=S_{1}(T,1,0)S_{2}(T+1,1,0)=S_{1}(T,0,1)S_{2}(T+1,0,1)$, for any $t\in\mathbb{R}$, $E(t,1)-aR(1,0)$ is not 00. ∎
Proposition 3.9.
When $l$ is odd and $n$ is at least $3$, the approximate solution $\frac{1}{a}E(T,X_{1})$ or $\frac{1}{a}E(T-X_{1},X_{1})$ associated with (*3) is a primitive approximate solution.
Proof.
Recall the equation $\frac{1}{a}E(t,1)-R(1,z)=\frac{1}{a}S_{1}(t,1,z)S_{2}(t+1,1,z)$. Note that $l$ is odd implies $l+1\geq 2n-1-l>1$. So by Lemma 3.3, we have $\frac{1}{a}e_{2n}=0$ or $\frac{1}{a}e_{2n}=1$. And when $\frac{1}{a}e_{2n}$ is 00, $\frac{1}{a}E(0,1)-R(1,\exp\frac{2\pi i}{2n+2})$ vanishes; when $\frac{1}{a}e_{2n}$ is $1$, $\frac{1}{a}E(0,1)-R(1,\exp\frac{2\pi i}{2n})$ vanishes.
Suppose that $n$ is $3$. When $\frac{1}{a}e_{2n}$ is 00, we have $\frac{1}{a}S_{1}(0,1,z)S_{2}(1,1,z)=-(z^{6}+z^{4}+z^{2}+1)$. The prime factorization of $z^{6}+z^{4}+z^{2}+1$ over $\mathbb{Q}[z]$ is $(z^{4}+1)(z^{2}+1)$. Note $\deg(S_{1})\geq\deg(S_{2})$ and $\deg(S_{2})>1$, we must have $S_{1}(0,1,z)=\lambda(z^{4}+1)$ for some $\lambda\in\mathbb{Q}$. Then $\exp\frac{2\pi i}{8}$ is a root of $S_{1}(0,1,z)$ and hence $\frac{1}{a}E(T,X_{1})$ is primitive. When $\frac{1}{a}e_{2n}$ is $1$, we have $\frac{1}{a}S_{1}(0,1,z)S_{2}(1,1,z)=-(z^{6}+z^{4}+z^{2})=-z^{2}(z^{4}+z^{2}+1)=-z^{2}(z^{2}+z+1)(z^{2}-z+1)$. Note that $S_{1}$ is symmetric in $X_{1},X_{2}$ and $\deg(S_{1})$ is at least $\deg(S_{2})$, we have $S_{1}(0,1,z)=\lambda(z^{4}+z^{2}+1)$. Then $\exp\frac{2\pi i}{6}$ is a root of $S_{1}(0,1,z)$ and hence $\frac{1}{a}E(T,X_{1})$ is also primitive.
Suppose now $n$ is at least $4$ and $\frac{1}{a}E(T,X_{1})$ is not a primitive solution. Then we have
$S_{2}(1,1,\exp\frac{2\pi i}{2n+2})=0\text{ or }S_{2}(1,1,\exp\frac{2\pi i}{2n})=0.$ (3.3)
We now show that $\frac{1}{a}E(T-X_{1},X_{1})$ is a primitive solution. First we have the following inequalities:
$\frac{\pi}{2n}<\frac{2\pi}{2n+2}<\frac{2\pi}{2n}<\frac{3\pi}{2n}<\frac{4\pi}{2n+2}.$
(Note that for the last inequality, we use the condition $n\geq 4$). Since the roots of $S_{2}(0,1,z)$ satisfy $z^{2n}=1$ or $z^{2n+2}=1$, it suffices to show that $S_{2}(0,1,z)$ has a non-zero root $y_{0}$ with argument satisfying $\frac{\pi}{2n}<\arg y_{0}<\frac{3\pi}{2n}$. By the above inequalities, we have $y_{0}=\frac{2\pi i}{2n+2}$ or $y_{0}=\frac{2\pi i}{2n}$. Then $\frac{1}{a}E(T-X_{1},X_{1})$ is a primitive approximate solution.
It is enough to show that for any $t\in\mathbb{R}$, $S_{2}(t+1,1,z)$ as a polynomial of $z$ has a non-zero root with argument in $(\frac{\pi}{2n},\frac{3\pi}{2n})$. Denote by $K$ the set $\{t\in\mathbb{R}~{}|~{}\exists~{}r(t)(\neq 0)\in\mathbb{C},~{}S_{2}(t+1,1,r(t))=0\text{ with }\frac{\pi}{2n}<\arg r(t)<\frac{3\pi}{2n}\}$. By (3.3), we have $0\in K$. Note that $K$ is open by construction, if we can show $K$ is closed, then $K$ is $\mathbb{R}$. Let $t_{0}$ be a limit point of $K$. Then $S_{2}(t_{0}+1,1,z)$ has a root $r(t_{0})$ which is a limit point of $\{r(t)~{}|~{}t\in K\}$. By Lemma 3.8, for any $t\in\mathbb{R}$, as a polynomial in $z,$ the roots of $\frac{1}{a}E(t,1)-R(1,z)$ are non-zero. So the roots of $S_{2}(t+1,1,z)$ are also non-zero. In particular, $r(t_{0})$ is not zero. Suppose that $t_{0}\notin K$, then we have $\arg r(t_{0})=\frac{m\pi}{2n}$, where $m$ is $1$ or $3$. We may assume $r(t_{0})=\rho\exp\frac{im\pi}{2n}$, where $\rho$ is a positive real number. Then $\frac{1}{a}E(t_{0},1)-R(1,\rho\exp\frac{im\pi}{2n})$ is 00. To get a contradiction, we show for any $d\in\mathbb{R}$, we have $R(1,\rho\exp\frac{im\pi}{2n})+d\neq 0$. If $R(1,\rho\exp\frac{im\pi}{2n})+d=0$, we have the following identities:
$(R(1,\rho\exp\frac{im\pi}{2n})+d)(\rho^{2}\exp\frac{2im\pi}{2n}-1)$
$=$ $R(1,\rho\exp\frac{im\pi}{2n})(\rho^{2}\exp\frac{2im\pi}{2n}-1)+d(\rho^{2}\exp\frac{2im\pi}{2n}-1)$
$=$ $\rho^{2n+2}\exp\frac{im(2n+2)\pi}{2n}-1+d\rho^{2}\exp\frac{2im\pi}{2n}-d$
$=$ $(d\rho^{2}-\rho^{2n+2})\exp\frac{im\pi}{n}-(d+1)=0.$
Since $n$ is bigger than $3$ and $m$ is $1$ or $3$, $\exp\frac{im\pi}{n}$ is not real. We must have $d+1=0$ and $d\rho^{2}-\rho^{2n+2}=0$. But it implies $d\rho^{2}-\rho^{2n+2}=-\rho^{2}-\rho^{2n+2}=0$, which is absurd. ∎
Now we complete the proof for the case $l$ is odd. By Proposition 3.5 and Proposition 3.9, we have the following possibilities (note that the coefficient of $T^{2n}$ in $E(T,X_{1})$ is $1$):
$E(T,X_{1})$ is $T^{2n}$ or $a\Sigma_{n}(\frac{T^{2}}{b_{1}},X_{1}^{2})$; $E(T-X_{1},X_{1})$ is $T^{2n}$ or $a\Sigma_{n}(\frac{T^{2}}{b_{2}},X_{1}^{2})$, where $b_{i}(\in\mathbb{Q})$ satisfy $b_{i}^{n}=a$$(i=1,2)$.
Proposition 3.10.
The above possibilities are all impossible.
Proof.
If $E(T,X_{1})$ is $T^{2n}$, $T^{2n}-aR(X_{1},X_{2})$ is $S_{1}(T,X_{1},X_{2})S_{2}(T+X_{1},X_{1},X_{2})$. However, we have
$X_{1}-\exp\frac{2\pi i}{2n+2}X_{2}\mid R(X_{1},X_{2}),~{}(X_{1}-\exp\frac{2\pi i}{2n+2}X_{2})^{2}\nmid R(X_{1},X_{2}),~{}X_{1}-\exp\frac{2\pi i}{2n+2}X_{2}\nmid 1.$
By Eisenstein’s criterion, $T^{2n}-aR(X_{1},X_{2})$ is irreducible considered as the polynomial in the variable $T$ with coefficients in $\mathbb{Q}[X_{1},X_{2}]$. This leads to a contradiction. Similar arguments can be applied to the case $E(T-X_{1},X_{1})$ is $T^{2n}$.
If $E(T,X_{1})$ is $b_{1}^{n}\Sigma_{n}(\frac{T^{2}}{b_{1}},X_{1}^{2})(=\Sigma_{n}(T^{2},b_{1}X_{1}^{2}))$, then we have the equalities $E(T,X_{1})-aR(X_{1},X_{2})=\Sigma_{n}(T^{2},b_{1}X_{1}^{2})-\Sigma_{n}(b_{1}X_{1}^{2},b_{1}X_{2}^{2})=(T^{2}-b_{1}X_{2}^{2})\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})$ (for the last equality, see [2, Section 7.2] for example). We will make use of the following claim, whose proof will be given in Proposition B.3.
Claim: $\Sigma_{n-1}(T^{2},X_{1}^{2},X_{2}^{2})$ is irreducible in $\mathbb{C}[T,X_{1},X_{2}]$.
As $b_{1}$ is not 00, $\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})$ is irreducible. We have $(T^{2}-b_{1}X_{2}^{2})\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})=S_{1}(T,X_{1},X_{2})S_{2}(T+X_{1},X_{1},X_{2})$. Since $S_{1}(T,X_{1},X_{2})$ is symmetric in $X_{1},X_{2}$ and the coefficient of $T^{2n}$ of $S_{1}$ is $1$, then $S_{1}(T,X_{1},X_{2})$ is $\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})$ and hence $S_{2}(T+X_{1},X_{1},X_{2})$ is $T^{2}-b_{1}X_{2}^{2}$. So $S_{2}(T,X_{1},X_{2})$ is $(T-X_{1})^{2}-b_{1}X_{2}^{2},$ which is not symmetric in $X_{1},X_{2}.$ We get a contradiction.
If $E(T-X_{1},X_{1})$ is $b_{2}^{n}\Sigma_{n}(\frac{T^{2}}{b_{2}},X_{1}^{2})(=\Sigma_{n}(T^{2},b_{2}X_{1}^{2}))$, we have $(T^{2}-b_{2}X_{2}^{2})\Sigma_{n-1}(T^{2},b_{2}X_{1}^{2},b_{2}X_{2}^{2})=S_{1}(T-X_{1},X_{1},X_{2})S_{2}(T,X_{1},X_{2})$. Similarly, $S_{2}(T,X_{1},X_{2})$ is $\Sigma_{n-1}(T^{2},b_{2}X_{1}^{2},b_{2}X_{2}^{2})$. But $\deg(S_{2})$ is at most $\deg(S_{1})$ and $n$ is at least $3$, it is impossible. ∎
Now we prove our main theorem.
Theorem 3.11.
Assume $n$ is at least $3$, then every uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}$ or $\mathbb{Q}^{2n+2}$ splits and $\mu(\mathbb{Q}^{2n+1})$ is $2n$.
Proof.
Combining Proposition 2.1, Proposition 3.6, Proposition 3.7 and Proposition 3.10, we can show that every uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}$ splits for $n\geq 3$. As the tangent bundle $T_{\mathbb{Q}^{2n+1}}$ is unsplit, the threshold $\mu(\mathbb{Q}^{2n+1})$ is $2n$.
Now let $E^{\prime}$ be a uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+2}$. For every smooth hyperplane section $\mathbb{Q}^{2n+1}\hookrightarrow\mathbb{Q}^{2n+2}$, the restriction $E^{\prime}|_{\mathbb{Q}^{2n+1}}$ is a uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}$, hence $E^{\prime}|_{\mathbb{Q}^{2n+1}}$ splits. So by [9, Corollary 3.3], $E^{\prime}$ splits. ∎
Remark 3.12.
For a majority of generalized Grassmannians $X$, $\mu(X)$ is $e.d.(\mathrm{VMRT})$ (see [4, Page 3, Table 2]). Note that the $e.d.(\mathrm{VMRT})$ of both $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}$ are $12n-$ 1. Theorem 3.11 shows that the splitting thresholds for uniform vector bundles on $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}(n\geq 3)$ are at least $2n$. So $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}(n\geq 3)$ are the first known examples such that $\mu(X)$ is bigger than $e.d.(\mathrm{VMRT})$.
In [4], the authors classify all unsplit uniform bundles of minimal rank on the generalized Grassmannians $B_{n}/P_{k}~{}(2\leq k<\frac{2n}{3})$, $B_{n}/P_{n-2}$, $B_{n}/P_{n-1}$ and $D_{n}/P_{k}~{}(2\leq k<\frac{2n-2}{3})$, $D_{n}/P_{n-3}$, $D_{n}/P_{n-2}$. As direct corollaries of Theorem 3.11, we can give further classification results for uniform bundles on $B_{n}/P_{k}~{}(\frac{2n}{3}\leq k\leq n-3)$ and $D_{n}/P_{k}~{}(\frac{2n-2}{3}\leq k\leq n-4)$.
Corollary 3.13.
Let $X$ be $B_{n}/P_{k}$, where $k$ is $\frac{2n}{3}$ and is at least $6$. Let $E$ be a uniform vector bundle on $X$ of rank $r$.
  • If $r$ is smaller than $k$, then $E$ is a direct sum of line bundles.
  • If $r$ is $k$, then $E$ is either a direct sum of line bundles or $E_{\lambda_{1}}\otimes L$ or $E_{\lambda_{1}}^{\vee}\otimes L$ for some line bundle $L$, where $E_{\lambda_{1}}$ is the irreducible homogeneous bundle corresponding to the highest weight $\lambda_{1}$.
Proof.
Note that $e.d.(\mathrm{VMRT})$ of $X$ is $k-1(=2n-2k-1)$. By [4, Theorem 1.1 (1)], the first assertion follows. For the case $r=k$, since $12n-2k+1=k+$ 1 is at least $7$, $E|_{\mathbb{Q}^{2n-2k+1}}$ splits by Theorem 3.11. Suppose $E$ is unsplit, then $E|_{\mathbb{P}^{k}}$ is also unsplit. The second assertion then follows from [4, Proposition 4.4]. ∎
Corollary 3.14.
Let $X$ be $B_{n}/P_{k}$ with $\frac{2n}{3}<k\leq n-3$. Every uniform bundle of rank $2n-2k$ on $X$ splits.
Proof.
If $E$ is a uniform bundle of rank $2n-2k$ on $X$, then $E|_{\mathbb{P}^{k}}$ splits, as $2n-2k$ is smaller than $k$. On the other hand, since $k$ is at most $n-3$ and hence $12n-2k+$ 1 is at least $7$, $E|_{\mathbb{Q}^{2n-2k+1}}$ splits by Theorem 3.11. Because any $2$-plane in $X$ is contained in a $\mathbb{P}^{k}$ or $\mathbb{Q}^{2n-2k+1}$, $E$ splits by [1, Corollary 3.6]. ∎
Similar to the proofs of the above corollaries, we can prove the following results.
Corollary 3.15.
Let $X$ be $D_{n}/P_{k}$, where $k$ is $\frac{2n-2}{3}$ and $k$ is at least $6$. Let $E$ be a uniform vector bundle on $X$ of rank $r$.
  • If $r$ is smaller than $k$, then $E$ is a direct sum of line bundles.
  • If $r$ is $k$, then $E$ is either a direct sum of line bundles or $E_{\lambda_{1}}\otimes L$ or $E_{\lambda_{1}}^{\vee}\otimes L$ for some line bundle $L$, where $E_{\lambda_{1}}$ is the irreducible homogeneous bundle corresponding to the highest weight $\lambda_{1}$.
Corollary 3.16.
Let $X$ be $D_{n}/P_{k}$ with $\frac{2n-2}{3}<k\leq n-4$. Every uniform bundle of rank $22n-2k-$ 2 on $X$ splits.
4. Splitting type of unsplit uniform bundle of minimal rank
There are some restrictions on the splitting types of unsplit uniform bundles of minimal rank. The following theorem generalizes [1, Corollary 4.7] to generalized Grassmannians associated with short roots.
Theorem 4.1.
Assume $E$ is an unsplit uniform bundle on a generalized Grassmannian $X$ whose rank is $\mu(X)+1$. If the splitting type of $E$ is $(a_{1},a_{2},\dots,a_{r})$ $(a_{1}\geq a_{2}\geq\cdots\geq a_{r})$, then $max\{a_{i}-a_{i+1}|1\leq i\leq r-1\}$ is $1$.
We now prove Theorem 4.1.
Proof.
Let $L$ be a line in $X$. We denote the inclusion morphism by $f_{L}:L(\simeq\mathbb{P}^{1})\rightarrow X$. As $X$ is a homogeneous variety, the tangent bundle $T_{X}$ is globally generated. Then $f_{L}^{*}(T_{X})$ is also globally generated and hence $H^{1}(L,f_{L}^{*}(T_{X}))$ vanishes. So Mor$(\mathbb{P}^{1},X)$ is smooth at $[f_{L}]$.
If $max\{a_{i}-a_{i+1}|1\leq i\leq r-1\}$ is 00, by [10, Theorem 1.2], $E$ is trivial. If $max\{a_{i}-a_{i+1}|1\leq i\leq r-1\}$ is at least $2$, there would exist a number $j<r$ such that $a_{j}-a_{j+1}\geq 2$. By [11, Proposition 3.1], there would be a subbundle $W$ of $E$ satisfying the following two properties:
  • $W$ is a uniform bundle of splitting type $(a_{1},\dots,a_{j})$.
  • the quotient $U\triangleq E/W$ is a uniform bundle of splitting type $(a_{j+1},\dots,a_{r})$.
As rk$(U)$ and rk$(W)$ are at most $\mu(X)$, both $U$ and $W$ split. Since $\text{Ext}^{1}(U,W)$ vanishes, the exact sequence $0\rightarrow W\rightarrow E\rightarrow U\rightarrow 0$ would split. So $E(\simeq U\oplus W)$ would be a direct sum of line bundles. ∎
In [3, Page 29, Conjecture], Ellia proposes a conjecture that every uniform bundle on $\mathbb{P}^{n}$ of splitting type
$(\underbrace{a_{1},\dots,a_{1}}_{l_{1}},\underbrace{a_{2},\dots,a_{2}}_{l_{2}},\dots,\underbrace{a_{k},\dots,a_{k}}_{l_{k}})$
with $a_{i}>a_{i+1}$ and $l_{i}\leq n-1$ for any $1\leq i\leq k$ necessarily splits. Using the same argument as in Theorem 4.1, we can reduce this conjecture to the case $l_{i}\leq n-1$ and $a_{i}-a_{i+1}=1$ for any $11\leq i\leq k-$ 1. In particular, the following Theorem 4.2 partially answers Ellia’s conjecture in a more general setting.
Theorem 4.2.
Given a generalized Grassmannian $X$. Let $E$ be a uniform bundle on $X$. Assume that the splitting type of $E$ is $(\underbrace{a_{1},\dots,a_{1}}_{l_{1}},\underbrace{a_{2},\dots,a_{2}}_{l_{2}},\dots,\underbrace{a_{k},\dots,a_{k}}_{l_{k}})$ with $a_{i}>a_{i+1}$. If for any $11\leq i\leq k-$ 1, $a_{i}-a_{i+1}$ is at least $2$, then $E$ splits.
Proof.
We prove by induction on $k$. The case $k=1$ is obviously true. Suppose the case $k<n$ is true. We now prove the case $k=n$.
Following the same strategy in the proof of Theorem 4.1, by [11, Proposition 3.1], there exists a subbundle $W$ of $E$ satisfying the following two properties:
  • $W$ is a uniform bundle of splitting type $(\underbrace{a_{1},\dots,a_{1}}_{l_{1}})$.
  • the quotient $U\triangleq E/W$ is a uniform bundle of splitting type $(\underbrace{a_{2},\dots,a_{2}}_{l_{2}},\dots,\underbrace{a_{n},\dots,a_{n}}_{l_{n}})$
Then $W$ is isomorphic to $\mathcal{O}_{X}^{\oplus l_{1}}(a_{1})$ where $\mathcal{O}_{X}(1)$ is the ample generator of Pic$(X)$. By induction, the bundle $U$ splits. Since $\text{Ext}^{1}(U,W)$ vanishes, the exact sequence $0\rightarrow W\rightarrow E\rightarrow U\rightarrow 0$ splits. So $E(\simeq U\oplus W)$ is a direct sum of line bundles.
Appendix A Additional details for Proposition 2.1
In this appendix, we use notations as in the proof of Proposition 2.1 and provide more details by sketching the proof of [14, Page 204, Case (ii)] and [13, Lemma 3.3 (ii) and Case 1 of (iii)].
Suppose that there exists a non-constant morphism $f:\mathbb{Q}^{2n+1}\rightarrow\mathbb{G}(l,2n+1)$. Let $g(t)$ and $h(t)$ be the polynomials in Proposition 2.1. We set
$g(t)=1-c_{1}t+c_{2}t^{2}+\cdots+(-1)^{l+1}c_{l+1}t^{l+1},$
$h(t)=1+d_{1}t+d_{2}t^{2}+\cdots+d_{2n+1-l}t^{2n+1-l}.$
Let $a$ be $2n+2]{c_{l+1}d_{2n+1-l}}$. Then, we have the equation (see Page 3, the equation (2.1))
$g(t)h(t)=1+(-1)^{l+1}c_{l+1}d_{2n+1-l}t^{2n+2}=1+(-1)^{l+1}a^{2n+2}t^{2n+2}.$ (A.1)
First, we use the same method of [13, Lemma 3.3] to prove that $a$ is a rational number.
Proposition A.1.
We have the identities $t^{l+1}g(\frac{1}{at})=(-1)^{l+1}g(\frac{t}{a})$ and $t^{2n+1-l}h(\frac{1}{at})=h(\frac{t}{a})$.
Proof.
Set $g(t)=(1-\alpha_{1}at)(1-\alpha_{2}at)\cdots(1-\alpha_{l+1}at)$. By formula (A.1), we have
$|\alpha_{i}|=1,~{}\alpha_{i}\neq\alpha_{j}\text{ if }i\neq j,~{}\alpha_{i}^{-1}=\overline{\alpha_{i}}\in\{\alpha_{1},\alpha_{2},\dots,\alpha_{l+1}\}.$ (A.2)
Furthermore, $\alpha_{1}\cdots\alpha_{l+1}a^{l+1}$ equals $c_{l+1}$. Since both $a$ and $c_{l+1}$ are positive real numbers, we must have $\alpha_{1}\cdots\alpha_{l+1}=1$. Combining with (A.2), we get
$t^{l+1}g(\frac{1}{at})$ $=(t-\alpha_{1})(t-\alpha_{2})\cdots(t-\alpha_{l+1})=(\alpha_{1}^{-1}t-1)(\alpha_{2}^{-1}t-1)\cdots(\alpha_{l+1}^{-1}t-1)\alpha_{1}\cdots\alpha_{l+1}$
$=(\alpha_{1}t-1)(\alpha_{2}t-1)\cdots(\alpha_{l+1}t-1)=(-1)^{l+1}g(\frac{t}{a}).$
Similarly, we set $h(t)=(1+\alpha_{1}^{\prime}at)(1+\alpha_{2}^{\prime}at)\cdots(1+\alpha_{2n+1-l}^{\prime}at)$ and apply similar arguments to obtain $t^{2n+1-l}h(\frac{1}{at})=h(\frac{t}{a})$. ∎
Corollary A.2.
The number $a$ is $\frac{c_{m+1}}{c_{m}}$, where $m$ is $\frac{l}{2}$.
Proof.
The equation $t^{l+1}g(\frac{1}{at})=(-1)^{l+1}g(\frac{t}{a})$ shows that $c_{i}a^{-i}$ equals $c_{l+1-i}a^{-l-1+i}$ for $1\leq i\leq l$. We get the desired conclusion by taking $i$ to be $\frac{l}{2}$. ∎
Once one proves that $a$ is rational, as in the proof of Proposition 2.1, both $a,C_{i}$ and $D_{j}$ are positive integers, we claim that
Claim: If $a,C_{i}$ and $D_{j}$ are all positive integers, we can apply the same proof of [14, Case (ii)] to get a contradiction.
Now we sketch the proof of [14, Page 204, Case (ii)]. Let $G(t)$ be $g(\frac{t}{a})(=1-C_{1}t+C_{2}t^{2}+\cdots+(-1)^{l+1}C_{l+1}t^{l+1})$ and $H(t)$ be $h(\frac{t}{a})(=1+D_{1}t+D_{2}t^{2}+\cdots+D_{2n+1-l}t^{2n+1-l})$. From formula (A.1), we have $H(t)G(t)=1+(-1)^{l+1}t^{2n+2}$.
Lemma A.3.
Suppose that there exists a non-constant morphism $f$ from $\mathbb{Q}^{2n+1}$ to $\mathbb{G}(l,2n+1)$. Then we have the equations $n=l$, $C_{1}=C_{2}=\cdots=C_{l}=D_{1}=D_{2}=\cdots=D_{2n-l}=2$ and $C_{l+1}=D_{2n+1-l}=1$.
Proof.
Since $\mathbb{G}(l,2n+1)$ is isomorphic to $\mathbb{G}(2n+1-l,2n+1)$, we may assume the inequality $n\geq l$. From the equation $H(t)G(t)=1+(-1)^{l+1}t^{2n+2}$ and the fact that $C_{i},D_{j}$ are positive integers, the first part of Tango’s proof allows us to conclude that $H(1)$ equals $22n+$ 2 and $D_{1}$ is bigger than $1$.
Next we prove that $D_{1},D_{2},\dots,D_{2n-l}$ are not less than $2$. Suppose that there exists a positive integer $k(\leq 2n-l)$ such that $D_{k}$ equals $1$. Let $r$ be $\min\{k|D_{k}=1\}$. Then both $D_{1}$ and $D_{r-1}$ are bigger than $1$. From $H(t)=t^{2n+1-l}H(\frac{1}{t})$ (Proposition A.1), we obtain $D_{2n+1-l}=1$ and $D_{2n+1-i}=D_{i}$ for $1\leq i\leq 2n-l$. In particular, $D_{2n+1-r}$ is $D_{r}$. By the definition of $r$, we have $2r\leq 2n+1-l$. Let $Q$ be the universal quotient bundle of $\mathbb{G}(l,2n+1)$. By Pieri’s formula, the class
$f^{*}\omega_{r,r,0,\dots,0}$ $=f^{*}(\omega_{r,0,\dots,0}^{2}-\omega_{r+1,0,\dots,0}\omega_{r-1,0,\dots,0})$
$=f^{*}(c_{r}(Q))^{2}-f^{*}(c_{r+1}(Q))f^{*}(c_{r-1}(Q))=(D_{r}^{2}-D_{r+1}D_{r-1})a^{2r}H^{2r}$
is the pullback of a Schubert cycle, which is numerically non-negative. (For the definition of the Schubert cycle $\omega_{a_{0},\dots,a_{l}}$, we refer to [14]). So $D_{r}^{2}-D_{r+1}D_{r-1}\geq 0$. From the inequalities $D_{r}^{2}-D_{r+1}D_{r-1}\leq 1-2<0$, we get a contradiction. Hence, we have $2n+2=H(1)=1+D_{1}+\cdots+D_{2n+1-l}\geq 1+2(2n-l)+1=2(2n+1-l)$. Combining with the assumption $n\geq l$, we must have $n=l$, $D_{1}=D_{2}=\cdots=D_{2n-l}=2$ and $D_{2n+1-l}=1$. Finally, from $H(t)G(t)=1+(-1)^{l+1}t^{2n+2}$, we also have $C_{1}=C_{2}=\cdots=C_{l}=2$ and $C_{l+1}=1$. ∎
Now we return to the proof of Claim. Since $l$ is even and $(l,2n+1)$ is not $(2,5)$, we have $l\geq 4$ and $D_{1}=D_{2}=D_{3}=2$. Similar as above, we have $f^{*}\omega_{2,2,0,\dots,0}=f^{*}(\omega_{2,0,\dots,0}^{2}-\omega_{3,0,\dots,0}\omega_{1,0,\dots,0})=(D_{2}^{2}-D_{3}D_{1})a^{4}H^{4}=0$. By Pieri’s formula again, it shows that
$0=f^{*}(\omega_{2,2,0,\dots,0}\omega_{2n-1-l,0,\dots,0})=f^{*}\omega_{2n+1-l,2,0,\dots,0}+f^{*}\omega_{2n-l,2,1,0,\dots,0}+f^{*}\omega_{2n-l-1,2,2,0,\dots,0}.$
Since all the classes in the right hand side is numerically non-negative, we have $f^{*}\omega_{2n+1-l,2,0,\dots,0}=0$. On the other hand, we have
$f^{*}\omega_{2n+1-l,2,0,\dots,0}=f^{*}\omega_{2n-1-l,0,\dots,0}f^{*}\omega_{2,0,\dots,0}$
$=$ $f^{*}(c_{2n+1-l}(Q))f^{*}(c_{2}(Q))=D_{2n+1-l}D_{2}a^{2n+3-l}H^{2n+3-l}\neq 0,$
which is a contradiction.
Appendix B Classification of primitive approximate solutions
We use methods in [2] to classify primitive approximate solutions. Let $P(T,X_{1})=\sum_{k=0}^{2n}p_{k}X_{1}^{k}T^{2n-k}$ be a primitive approximate solution.
Proposition B.1.
If $p_{2n}$ is 00, then $P(T,X_{1})$ is $bT^{2n}$ for some rational number $b$.
Proof.
Let $S_{0}(T,X_{1},X_{2})$ be a symmetric divisor of $P(T,X_{1})-R(X_{1},X_{2})$ such that for a $2(n+1)$-th primitive unit root $y_{0}$, we have $S_{0}(0,1,y_{0})=0$. Since $y_{0}$ is a simple root of $P(0,1)-R(1,z)$, $y_{0}$ is also a simple root of $S_{0}(0,1,z)$. Therefore, by the implicit function theorem, there is a germ of holomorphic function $y(x)$ in a neighborhood of $x=0$ satisfying
$S_{0}(x(1+y(x)),1,y(x))=0\text{ and }y(0)=y_{0}.$
We now show by induction that for $m=1,\dots,2n-1$, we have $y^{(m)}(0)=p_{2n-m}=0$. As $S_{0}$ is symmetric in $X_{1}$ and $X_{2}$, we have
$P(x(1+y(x)),1)-R(y(x),1)=0,$ (1)
$P(x(1+y(x)),y(x))-R(y(x),1)=0.$ (2)
By taking the derivatives of (1) and (2) at $x=0$ and noting that $p_{2n}$ is 00, we obtain
$p_{2n-1}(1+y_{0})-R^{\prime}(y_{0},1)y^{\prime}(0)=0,$ (1’)
$p_{2n-1}(1+y_{0})y_{0}^{2n-1}-R^{\prime}(y_{0},1)y^{\prime}(0)=0.$ (2’)
From $R(y,1)(y^{2}-1)=y^{2n+2}-1$, we have $R^{\prime}(y,1)(y^{2}-1)+R(y,1)\cdot(2y)=(2n+2)y^{2n+1}$. As $R(y_{0},1)$ is 00 and $y_{0}$ is not ±$\pm 1$, $R^{\prime}(y_{0},1)$ is $\frac{(2n+2)y_{0}^{2n+1}}{y_{0}^{2}-1}(\neq 0)$. From (1’) and (2’), we get $p_{2n-1}(1+y_{0})(y_{0}^{2n-1}-1)=0$. Since $y_{0}$ is primitive, $y_{0}^{2n-1}-1$ and $1+y_{0}$ are not zero. We have $p_{2n-1}=0$. By (1’) and noting that $R^{\prime}(y_{0},1)\neq 0$, $y^{\prime}(0)$ is 00.
For the case $m\geq 2$, we assume by induction that $y^{(m^{\prime})}(0)=p_{2n-m^{\prime}}=0$ for $m^{\prime}<m$. Then the $m$-th derivatives of (1) and (2) satisfy the following equations:
$m!p_{2n-m}(1+y_{0})^{m}-R^{\prime}(y_{0},1)y^{(m)}(0)=0,$ ($1^{m}$)
$m!p_{2n-m}(1+y_{0})^{m}y_{0}^{2n-m}-R^{\prime}(y_{0},1)y^{(m)}(0)=0.$ ($2^{m}$)
Since $y_{0}$ is primitive, $y_{0}^{2n-m}-1$ is not zero. We have $y^{(m)}(0)=p_{2n-m}=0$ as above. ∎
Proposition B.2.
If $p_{2n}$ is $1$, then $P(T,X_{1})$ is $\Sigma_{n}(bT^{2},X_{1}^{2})$ for some $b\in\mathbb{Q}$.
Proof.
Let $S_{+}(T,X_{1},X_{2})$ be a symmetric divisor of $P(T,X_{1})-R(X_{1},X_{2})$ such that for a $2n$-th primitive unit root $y_{0}$, we have $S_{+}(0,1,y_{0})=0$. Note the equation $\Sigma_{n}(p_{2n-2}T^{2},X_{1}^{2})-R(X_{1},X_{2})=(p_{2n-2}T^{2}-X_{2}^{2})\Sigma_{n-1}(p_{2n-2}T^{2},X_{1}^{2},X_{2}^{2})$ (see [2, Section 7.2] for example). Let $S_{-}(T,X_{1},X_{2})$ be $\Sigma_{n-1}(p_{2n-2}T^{2},X_{1}^{2},X_{2}^{2})$, then $S_{-}(0,1,y_{0})$ is also 00. Denote $P(T,X_{1})$ by $P_{+}(T,X_{1})$ and denote $\Sigma_{n}(p_{2n-2}T^{2},X_{1}^{2})$ by $P_{-}(T,X_{1})$. We are going to show that $P_{+}$ equals $P_{-}$. Let $y_{\pm}(x)$ be germs of holomorphic functions satisfying
$S_{\pm}(x(1+y_{\pm}(x)),1,y_{\pm}(x))=0\text{ and }y_{\pm}(0)=y_{0}.$
As $S_{\pm}$ is symmetric in $X_{1}$ and $X_{2}$, we have equations:
$P_{\pm}(x(1+y_{\pm}(x)),1)-R(y_{\pm}(x),1)=0,$ ($1_{\pm}$)
$P_{\pm}(x(1+y_{\pm}(x)),y_{\pm}(x))-R(y_{\pm}(x),1)=0.$ ($2_{\pm}$)
Let $p^{+}_{2n-m}$ (resp. $p^{-}_{2n-m}$) be the coefficient of $X_{1}^{2n-m}T^{m}$ in $P_{+}(T,X_{1})$ (resp. $P_{-}(T,X_{1})$). We prove $p^{+}_{2n-m}=p^{-}_{2n-m}$ and $y^{(m)}_{+}(0)=y^{(m)}_{-}(0)$ by induction for $m$ $(0\leq m\leq 2n)$.
When $m$ is 00, we have $p_{2n}^{+}=p_{2n}^{-}=1$, $y_{+}(0)=y_{-}(0)=y_{0}$. We next show $p_{2n-1}^{+}=p_{2n-1}^{-}=0$ and $y^{\prime}_{+}(0)=y^{\prime}_{-}(0)=0$. Taking the derivatives of ($1_{\pm}$) and ($2_{\pm}$) at $x=0$, we get
$p_{2n-1}^{\pm}(1+y_{0})-R^{\prime}(y_{0},1)y^{\prime}_{\pm}(0)=0,$ ($1^{\prime}_{\pm}$)
$p_{2n-1}^{\pm}(1+y_{0})y_{0}^{2n-1}+2ny_{0}^{2n-1}y^{\prime}_{\pm}(0)-R^{\prime}(y_{0},1)y^{\prime}_{\pm}(0)=0.$ ($2^{\prime}_{\pm}$)
From $R(y,1)(y^{2}-1)=y^{2n+2}-1$, we have $R^{\prime}(y,1)(y^{2}-1)+R(y,1)\cdot(2y)=(2n+2)y^{2n+1}$. Note $R(y_{0},1)=1$ (as $y_{0}$ is a $2n$-th primitive unit root). Substituting it in the above equation, we have
$2ny_{0}-y_{0}^{2}R^{\prime}(y_{0},1)=-R^{\prime}(y_{0},1).$ (†)
Multiplying ($2^{\prime}_{\pm}$) by $y_{0}^{2}$ and using the relation (B), one has
$p_{2n-1}^{\pm}(1+y_{0})y_{0}-R^{\prime}(y_{0},1)y_{\pm}^{\prime}(0)=0.$ ($y_{0}^{2}\cdot 2^{\prime}_{\pm}$)
$p_{2n-1}^{\pm}(1+y_{0})-R^{\prime}(y_{0},1)y_{\pm}^{\prime}(0)=0,$ ($1^{\prime}_{\pm}$)
Since $y_{0}$ is primitive, the number $\left|\begin{array}[]{ccc}y_{0}(1+y_{0})&-R^{\prime}(y_{0},1)\\ 1+y_{0}&-R^{\prime}(y_{0},1)\end{array}\right|=(1-y_{0})(1+y_{0})R^{\prime}(y_{0},1)$ is not 00. Then one obtains $p_{2n-1}^{\pm}=y_{\pm}^{\prime}(0)=0$.
When $m$ is at least $2$, by induction, we may assume $p^{+}_{2n-m^{\prime}}=p^{-}_{2n-m^{\prime}}$ and $y_{+}^{(m^{\prime})}(0)=y_{-}^{(m^{\prime})}(0)$ for $m^{\prime}<m$. By taking the $m$-th derivatives of ($1_{\pm}$) and ($2_{\pm}$) at $x=0$, we have
$m!p_{2n-m}^{\pm}(1+y_{0})^{m}-R^{\prime}(y_{0},1)y^{(m)}_{\pm}(0)+T^{1}_{\pm}=0,$ ($1^{m}_{\pm}$)
$m!p_{2n-m}^{\pm}(1+y_{0})^{m}y_{0}^{2n-m}+2ny_{0}^{2n-1}y^{(m)}_{\pm}(0)-R^{\prime}(y_{0},1)y^{(m)}_{\pm}(0)+T^{2}_{\pm}=0,$ ($2^{m}_{\pm}$)
where $T^{i}_{\pm}$ are the remaining terms satisfying $T^{1}_{-}=T^{1}_{+}$ and $T^{2}_{-}=T^{2}_{+}$. For the case $m=2$, by construction, we automatically have $p_{2n-2}^{-}=p_{2n-2}=p_{2n-2}^{+}$. Then one obtains $y_{+}^{(2)}(0)=y_{-}^{(2)}(0)$ from ($1^{m}_{\pm}$). Suppose now $m$ is at least $3$, multiplying ($2^{m}_{\pm}$) by $y_{0}^{2}$ and using (B), we have
$m!p_{2n-m}^{\pm}(1+y_{0})^{m}y_{0}^{2n-m+2}-R^{\prime}(y_{0},1)y_{\pm}^{(m)}(0)+y_{0}^{2}T^{2}_{\pm}=0.$ ($y_{0}^{2}\cdot 2^{m}_{\pm}$)
Note that $\left|\begin{array}[]{ccc}m!(1+y_{0})^{m}y_{0}^{2n-m+2}&-R^{\prime}(y_{0},1)\\ m!(1+y_{0})^{m}&-R^{\prime}(y_{0},1)\end{array}\right|=m!(1-y_{0}^{2n-m+2})(1+y_{0})^{m}R^{\prime}(y_{0},1)$ is not zero for $m\geq 3$. By solving the system of linear equations {($y_{0}^{2}\cdot 2^{m}_{\pm}$), ($1^{m}_{\pm}$)} (view $p_{2n-m}^{\pm}$ and $y_{\pm}^{(m)}(0)$ as indeterminate), we have $p_{2n-m}^{+}=p_{2n-m}^{-}$ and $y_{-}^{(m)}(0)=y_{+}^{(m)}(0)$. ∎
Proposition B.3.
$\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})$ is irreducible in $\mathbb{C}[T,X_{1},X_{2}]$.
Proof.
It suffices to show that the variety $V:=\{(T,X_{1},X_{2})|\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})=0\}\subset\mathbb{C}^{3}$ is smooth on $\mathbb{C}^{3}\backslash\{0\}$. By [2, Lemma 7.2], the variety defined by $\Sigma_{n}(T,X_{1},X_{2})=0$ is smooth on $\mathbb{C}^{3}\backslash 0$. Note that the map $(T,X_{1},X_{2})\mapsto(T^{2},X_{1}^{2},X_{2}^{2})$ is a local isomorphism outside the locus defined by $TX_{1}X_{2}=0$, $V$ is smooth on $\mathbb{C}^{3}\backslash\{TX_{1}X_{2}=0\}$.
Now suppose $(t,u,v)(\neq(0,0,0))$ is a singular point of $V$, then one of $t,u$ and $v$ is 00. By symmetry, we assume that $t$ is 00. Then there are equations
$\frac{\partial}{\partial T}\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})|_{(0,u,v)}=\frac{\partial}{\partial X_{1}}\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})|_{(0,u,v)}=\frac{\partial}{\partial X_{2}}\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})|_{(0,u,v)}=0.$
Furthermore, we have equations:
$\frac{\partial}{\partial X_{1}}\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})|_{(0,u,v)}=\frac{\partial}{\partial X_{1}}\Sigma_{n}(X_{1}^{2},X_{2}^{2})|_{(u,v)}=0,$
$\frac{\partial}{\partial X_{2}}\Sigma_{n}(T^{2},X_{1}^{2},X_{2}^{2})|_{(0,u,v)}=\frac{\partial}{\partial X_{2}}\Sigma_{n}(X_{1}^{2},X_{2}^{2})|_{(u,v)}=0.$
Without loss of generality, we may assume $v$ is not 00. As $\Sigma_{n}(u^{2},v^{2})(=\Sigma_{n}(0,u^{2},v^{2}))$ vanishes, $u/v$ is a root of $\Sigma_{n}(z^{2},1)=0$ with multiplicity greater than $2$. But $\Sigma_{n}(z^{2},1)=\frac{z^{2n+2}-1}{z^{2}-1}$ has no multiple roots, which is a contradiction. ∎